### v1

# Novel low-energy peak in the one-particle spectral function of the electron gas: Breakdown of the Landau’s hypothesis as to the Fermi liquids in simple metals

Abstract
Based on a nonperturbative scheme to determine the self-energy $\Sigma({\bm{k}},i\omega_{n})$ with automatically satisfying the Ward identity and the total momentum conservation law, a fully self-consistent calculation is done in the electron gas at various temperatures $T$ to obtain $G({\bm{k}},i\omega_{n})$ the one-particle Green’s function with fulfilling all known conservation laws, sum rules, and correct asymptotic behaviors; here, $T$ is taken unprecedentedly low, namely, $k_{\rm B}T/\varepsilon_{\rm F}$ down to $10^{-4}$ with $\varepsilon_{\rm F}$ the Fermi energy, and tiny mesh $\Delta k$ as small as $10^{-4}k_{\rm F}$ is chosen near the Fermi surface in ${\bm{k}}$ space with $k_{\rm F}$ the Fermi momentum. By analytically continuing $G({\bm{k}},i\omega_{n})$ to the retarded function $G^{R}({\bm{k}},\omega)$, we find a novel low-energy peak, in addition to the quasiparticle (QP) peak and one- and two-plasmon high-energy satellites, in the spectral function $A({\bm{k}},\omega)[=\!-{\rm Im}G^{R}({\bm{k}},\omega)/\pi]$ for $k_{\rm B}T\!\lesssim\!10^{-3}\varepsilon_{\rm F}$ in the simple-metal density region ($62\!<\!r_{s}\!<\!$ 6 with $r_{s}$ the dimensionless density parameter). This new peak is attributed to the effect of excitonic attraction on $\Sigma({\bm{k}},i\omega_{n})$ arising from multiple excitations of tightly-bound electron-hole pairs in the polarization function $\Pi({\bm{q}},i\omega_{q})$ for $|{\bm{q}}|\!\approx\!2k_{\rm F}$ and $|\omega_{q}|\!\ll\!\varepsilon_{\rm F}$ and thus it is dubbed “excitron”. Although this excitron peak height is only about a one-hundredth of that of QP, its excitation energy is about a half of that of QP for $|{\bm{k}}|\!\approx\!k_{\rm F}$, seemingly in contradiction to the Landau’s hypothesis as to the one-to-one correspondence of low-energy excitations between a free Fermi gas and an interacting normal Fermi liquid. As for the QP properties, our results of both the effective mass $m^{*}$ and the renormalization factor $z^{*}$ are in good agreement with those provided by recent quantum Monte Carlo simulations and available experiments.
I Introduction
The Landau’s phenomenological theory is very useful in explaining various common low-temperature properties of normal Fermi liquids [1, 2]. In particular, the class of simple metals is regarded as an archetype system to which this Landau Fermi-liquid theory (FLT) successfully applies. In FLT, low-lying excited states in a free Fermi gas are assumed to evolve smoothly to those of a normal Fermi liquid by adiabatically turning on mutual interactions, indicating a one-to-one correspondence between a bare particle specified by the dispersion $\varepsilon_{\bm{k}}$ in the free Fermi gas and a quasiparticle (a key concept in FLT) by $\tilde{\varepsilon}_{\bm{k}}$ and a long lifetime ${\tau}_{\bm{k}}$ in the Fermi liquid.
In terms of the one-particle spectral function $A({\bm{k}},\omega)$, a quantity to be observed by angle-resolved photoemission spectroscopy, this assumption of one-to-one correspondence amounts to the change of $A({\bm{k}},\omega)$ for $|{\bm{k}}|\!\approx\!k_{\rm F}$ ($k_{\rm F}$: the Fermi momentum) and $\omega\!\ll\!\varepsilon_{\rm F}$ ($\varepsilon_{\rm F}$: the Fermi energy) from $\delta(\omega\!-\!\varepsilon_{\bm{k}})$ in the free Fermi gas into a single sharp peak (with the width ${\tau}_{\bm{k}}^{-1}$) centered at $\omega\!=\!\tilde{\varepsilon}_{\bm{k}}$ plus a smooth incoherent background in the normal Fermi liquid. Hereafter we employ units in which $\hbar\!=\!k_{\rm B}\!=\!1$.
In accordance with this $A({\bm{k}},\omega)$, the thermal one-particle Green’s function $G(K)$ with $K\!\equiv\!\{{\bm{k}},i\omega_{n}\}$, a combined notation of momentum ${\bm{k}}$ and fermion Matsubara frequency $\omega_{n}$, is written for $|{\bm{k}}|\!\approx\!k_{\rm F}$ and $|\omega_{n}|\!\ll\!\varepsilon_{\rm F}$ as
$G(K)\!=\!\frac{z_{\bm{k}}}{i\omega_{n}\!-\!\tilde{\varepsilon}_{\bm{k}}\!+\!i\ {\rm sgn}(\omega_{n})(2\tau_{\bm{k}})^{-1}}\!+\!G_{\rm incoh}(K),$ (1)
where $z_{\bm{k}}$ is the quasiparticle renormalization factor and $G_{\rm incoh}(K)$ is a smooth function, corresponding to the incoherent background in $A({\bm{k}},\omega)$. As long as $z_{\bm{k}}\!\neq\!0$, the quasiparticle appears as a pole in $G(K)$. This pole-singularity is intimately connected with the condition that the self-energy $\Sigma(K)$ is smooth enough to be analytically expanded around $|{\bm{k}}|\!=\!k_{\rm F}$ and $\omega_{k}\!=\!0$. This quasiparticle concept is verified to infinite order in perturbation expansion in the quantum field theory [3, 4, 5, 6]. It is also confirmed in both the renormalization-group approach [7] and the multidimensional bosonization [8, 9, 10].
In the last several decades, we have witnessed that FLT breaks down in a number of interesting metals broadly referred to as non-Fermi liquids (NFLs). Examples include the one-dimensional (1D) Luttinger liquids [11, 12, 13, 14, 15], the so-called strange-metal phase in high-$T_{c}$ cuprates [16, 17, 18, 19, 20, 21], and the fluctuation regime around a quantum critical point (QCP) [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. Various routes to NFLs from normal Fermi liquids have been critically reviewed in several papers [36, 37, 38, 39, 40, 41, 42].
Guided by those review papers, we envisage a couple of routes to the breakdown of FLT in simple metals. The first one is related to an issue of the long-range Coulomb interaction $V({\bm{q}})$($=\!4\pi e^{2}/{\bm{q}}^{2}$). In fact, this has been investigated in detail in the past [43, 44, 45] and it is concluded that $V({\bm{q}})$ at $|{\bm{q}}|\!\to\!0$ is not singular enough to break FLT in 3D metals due to the screening effect. Incidentally, transverse electromagnetic fields give rise to unscreened long-range interactions, leading to the breakdown of FLT even in alkali metals [46, 47, 48], but this will not be observed experimentally; it occurs only at extremely low temperatures because of its very weak coupling controlled by the small factor $(k_{\rm F}/mc)^{2}$, where $m$ is the mass of a free electron and $c$ is the velocity of light.
The second route is concerned with the response function at $|{\bm{q}}|\!\approx\!2k_{\rm F}$ and possible $2k_{\rm F}$ singularities in it [49, 50, 9, 51, 45, 10, 52, 26, 53]. This is a problem which has not been examined in detail, especially in the context of 3D simple metals. The density response at low energies and short wavelengths will be an essential ingredient in the present study, as we now explain.
The low-lying excited states in simple metals are well described by those of a homogeneous 3D electron gas, an assembly of $N$ electrons embedded in a uniform positive rigid background. In a recent paper [54], it is shown in the low-density electron gas that an excitonic collective mode $\omega_{\rm ex}(q)$ made of correlated electron-hole pair excitations appears as a soft mode for $q\!\approx\!2k_{\rm F}$. If this $\omega_{\rm ex}(q)$ vanishes at $q\!=\!q_{c}\!\approx\!2k_{\rm F}$, a macroscopic number of electron-hole pairs are produced spontaneously to form a CDW state with the wave number $q_{c}$ which is exactly the state predicted by Overhauser [55]. In actual simple metals for which the density parameter $r_{s}$ defined by
$r_{s}=\frac{1}{\alpha k_{\rm F}a_{\rm B}}\ {\rm with}\ \alpha=\left(\frac{4}{9\pi}\right)^{1/3}\approx 0.5211$ (2)
and $a_{\rm B}$ the Bohr radius is in the range from 2 to 6, this kind of CDW does not seem to exist, but from the perspective of QCP physics, the correlated electron-hole pair excitations around $2k_{\rm F}$ may be regarded as quantum fluctuations around the quantum critical CDW transition point. In this regard, those bosonic excitations in simple metals can be considered as marginally relevant processes in the sense of renormalization group to break FLT.
In fact, the above-mentioned $\omega_{\rm ex}(q)$ mode is confirmed as an incipient excitonic mode in the dynamic structure factor $S({\bf q},\omega)$ at $|{\bm{q}}|\!\approx\!2k_{\rm F}$ by recent ab initio path integral Monte Carlo simulations performed in the electron gas for $r_{s}\!=\!2-10$ [56, 57, 58] and $r_{s}\!=\!2-36$ [59], where the mode is referred to as a “roton”.
In this paper, by exploiting the precise polarization function in the charge channel derived from Monte Carlo data for the density response, we will accurately calculate $\Sigma(K)$ in the 3D homogeneous electron gas for $r_{s}\!<\!6$ with the motivation to inspect the validity of the Landau FLT in simple metals. For this inspection, we are not allowed to apply a theory based on any kinds of perturbation expansion to this problem, because NFL cannot be described by perturbation expansion starting from the noninteracting Green’s function $G_{0}(K)$; actually, the so-called $GW$ approximation to the Hedin’s closed set of equations, derived by using the screened interaction $W$ as an expansion parameter [60, 61, 62, 63, 64, 65], and its refinements [66, 67, 68, 69] have not shown any indication of the breakdown of FLT. The same is true both in the many-body perturbation theory (MBPT) with using some appropriate local-field factor [70, 71, 72, 73, 74, 75, 76] and in our old studies in the effective-potential expansion (EPX) method [77, 78].
[…]
By applying this newly developed code to the electron gas at $r_{s}\!=\!2,07$, $3.25$, $3.93$, $4.86$, $5.20$, and $5.62$ relevant to Al, Li, Na, K, Rb, and Cs, respectively, we successfully obtain a convergent result of $\Sigma(K)$ at each $r_{s}$ in $10-100$ times iterative calculations, starting with $\Sigma_{0}(K)$ the self-energy in the random-phase approximation (RPA) (or the $G_{0}W_{0}$ approximation [87]). For $T\lesssim 10^{-3}\varepsilon_{\rm F}$, the obtained $\Sigma(K)$ is not smooth enough at $|{\bm{k}}|\!\approx\!k_{\rm F}$ and $\omega_{k}\!\approx\!0$ to safely confirm FLT; by a heuristic analysis, this anomalous behavior is found to be well described by a branch-cut singularity in $G(K)$, seemingly exhibiting a symptom of the possible breakdown of FLT.
Figure 1: Single-particle spectral function $A({\bm{k}},\omega)$ for the 3D homogeneous electron gas at $r_{s}\!=\!4.86$ corresponding to potassium, $T\!=\!0.001\varepsilon_{\rm F}$, and $|{\bm{k}}|\!=\!0.82k_{\rm F}$.
This situation is more clearly seen in the one-particle spectral function $A({\bm{k}},\omega)\!\equiv\!-{\rm Im}G^{R}({\bm{k}},\omega+i\gamma)/\pi$ with $\gamma\!=\!\pi T$ and $G^{R}({\bm{k}},\omega)$ the retarded Green’s function obtained by an analytic continuation of $G(K)$ in complex $\omega$ plane through Padè approximants. A typical example of $A({\bm{k}},\omega)$ is shown in Fig. 1 in which a new peak indicated by “excitron” appears in addition to the predominant QP peak and one- and two-plasmon satellites. It is evident from this figure that an energy resolution in angle-resolved photoemission spectroscopy (ARPES) should be small, i.e., of the order of 1 meV or less, so as to experimentally distinguish this weak excitron peak from the very strong QP peak. In the previous ARPES studies, on the other hand, the resolution was about 0.2-0.4eV in 1980s [88, 89, 90, 91], 80-200meV in 1990s [92], and still 30meV in 2020s [93]. This is probably the reason why the excitron peak has not been detected so far, though some interesting unresolved features were observed near the Fermi level in the past [88, 89, 91].
Although the excitron emerges, its spectral strength is quite weak and $G(K)$ is dominated by the QP pole singularity. Thus, the bulk properties of the electron gas in the simple-metal density region will be well explained in terms of FLT in which the quantitative evaluation of the QP effective mass $m^{*}$ is a long-standing yet hot issue. We can determine $m^{*}$ from the QP peak position at $|{\bm{k}}|$ in the vicinity of $k_{\rm F}$; our values for $m^{*}$ are in very good agreement with those given in recent quantum Monte Carlo simulations [84, 86], but they definitely deviate from those in diffusion Monte Carlo (DMC) method [94]. Note that DMC was feared to be an unreliable method in determining $m^{*}$ [95].
The same is true for the QP renormalization factor $z^{*}$; our values for $z^{*}$ determined by the jump of the momentum distribution function $n({\bm{k}})$ at $|{\bm{k}}|\!=\!k_{\rm F}$ are in good agreement with those in recent quantum Monte Carlo simulations [96, 97] and available experiments  [97, 98, 99].
The two-plasmon satellite is another challenging issue in the theoretical studies of photoemission in the electron gas in connection with experiments in simple metals. In the usual $GW$ and related schemes, it is known to be very difficult to provide this two-plasmon satellite structure, which urged Aryasetiawan et al. to invent a $GW$ plus cumulant-expansion approach [100], a method manually including the multi-plasmon satellites. Afterwards, many other works followed in that direction [101, 102, 103, 104, 105], whereas Pavlyukh et al. succeeded in obtaining the structure without resort to the cumulant expansion for the first time [106]. As seen in Fig. 1, our present method is the second one to accomplish the goal of obtaining the two-plasmon satellites without manually including the multi-plasmon satellites.
This paper is organized as follows: In Sec. II we explain our theoretical framework to accurately determine the self-energy in the 3D homogeneous electron gas through a nonperturbative iteration loop with rigorously satisfying the Ward identity and the total momentum conservation law. A proposed functional form for the vertex function $\Gamma(K,K\!+\!Q)$ in Sec. II.5 is one of the main results of this paper. In Sec. III we give the calculated results on various important properties in the electron gas, including $n({\bm{k}})$, $z^{*}$, and $m^{*}$. Among others, we find the new low-energy peak (dubbed “excitron”) and two-plasmon satellites in $A({\bm{k}},\omega)$. Although we have treated the electron densities corresponding to all simple metals, the main stress is laid on Na, because Na is known to be an almost perfect realization of the electron gas with $r_{s}=3.93$. In Sec. IV we make a detailed account of the excitron, ascribing it to the effect of excitonic attraction arising from the multiple excitations of tightly-bound electron-hole pairs. Finally in Sec. V we give summary of this paper and discussions on related problems and future issues. Appendixes A--E provide some details of our numerical calculation.
II Self-Energy Calculation Scheme
II.1 Hamiltonian
The Hamiltonian $H$ for a 3D homogeneous electron gas is written in second quantization as
$H=$ $\sum_{{\bm{k}}\sigma}\varepsilon_{{\bm{k}}}c_{{\bm{k}}\sigma}^{+}c_{{\bm{k}}\sigma}$
$+\frac{1}{2}\sum_{{\bm{q}}\neq{\bm{0}}}\sum_{{\bm{k}}\sigma}\sum_{{\bm{k}^{\prime}}\sigma^{\prime}}V({\bm{q}})c_{{\bm{k}}+{\bm{q}}\sigma}^{+}c_{{\bm{k}^{\prime}}-{\bm{q}}\sigma^{\prime}}^{+}c_{{\bm{k}^{\prime}}\sigma^{\prime}}c_{{\bm{k}}\sigma},$ (3)
where $c_{{\bm{k}}\sigma}$ is the annihilation operator of an electron with momentum ${\bm{k}}$ and spin $\sigma$ whose single-particle energy is given by $\varepsilon_{{\bm{k}}}\!\equiv\!{\bm{k}}^{2}/(2m)\!-\!\varepsilon_{\rm F}$ with $\varepsilon_{\rm F}\!=\!k_{\rm F}^{2}/(2m)$. We will consider the system in unit volume and the total number of electrons $N$ is nothing but the electron density $n$, given in terms of the Fermi momentum $k_{\rm F}$ as $n\!=\!k_{\rm F}^{3}/(3\pi^{2})$.
In this system, the correlation energy per electron $\varepsilon_{c}$ is already known accurately as a function of $r_{s}$ by the Green’s-Function Monte Carlo (GFMC) method [107] and the interpolation formulas to reproduce the GFMC data [108, 109]. With the use of $\varepsilon_{c}(r_{s})$, $\mu_{c}$ the correlation contribution to the chemical potential $\mu$ and the compressibility $\kappa$ are, respectively, obtained as
$\frac{\mu_{c}}{\varepsilon_{\rm F}}$ $=\alpha^{2}r_{s}^{2}\left(\varepsilon_{c}-\frac{r_{s}}{3}\varepsilon^{\prime}\right),$ (4)
$\frac{\kappa_{\rm F}}{\kappa}$ $=\frac{d\,\mu}{d\,\varepsilon_{\rm F}}=1-\frac{\alpha r_{s}}{\pi}+\frac{\alpha^{2}r_{s}^{3}}{6}\left(r_{s}\varepsilon_{c}^{\prime\prime}-2\varepsilon_{c}^{\prime}\right),$ (5)
where $\kappa$ is given through the thermodynamic relation of $\kappa\!=\!(d\,n/d\,\mu)/n^{2}$ and $\kappa_{\rm F}$ is the compressibility in the noninteracting electron gas, given by $\kappa_{\rm F}\!=\!D_{\rm F}/n^{2}\!=\!3m/(k_{\rm F}^{2}n)$ with $D_{\rm F}\!=\!d\,n/d\,\varepsilon_{\rm F}\!=\!mk_{\rm F}/\pi^{2}$ the density of states at the Fermi level in the noninteracting system.
II.2 One-particle Green’s function
The Dyson equation relates the one-particle Green’s function $G(K)$ with the self-energy $\Sigma(K)$ through
$G^{-1}(K)\!=\!i\omega_{n}\!+\!\mu_{x}\!+\!\mu_{c}\!-\!\varepsilon_{{\bm{k}}}-\Sigma(K).$ (6)
Here, $\mu$ is divided into $\mu\!=\!\varepsilon_{\rm F}+\mu_{x}+\mu_{c}$ with $\mu_{x}$ the exchange contribution to $\mu$, given by
$\frac{\mu_{x}}{\varepsilon_{\rm F}}=-\frac{2\alpha r_{s}}{\pi}.$ (7)
Let us divide $\Sigma(K)$ into odd and even parts in $\omega_{n}$ as
$\Sigma(K)=\left[1-Z(K)\right]i\omega_{n}\!+\!\chi(K),$ (8)
where both $Z(K)$ and $\chi(K)$ are not only even functions in $\omega_{n}$ but also real functions, as seen by combining Eq. (8) with the analyticity property $\Sigma({\bm{k}},-i\omega_{n})=\Sigma^{*}({\bm{k}},i\omega_{n})$. Then, $G^{-1}(K)$ is rewritten as
$G^{-1}(K)=Z(K)i\omega_{n}-E(K),$ (9)
with the introduction of $E(K)\!\equiv\!{\bm{k}}^{2}/(2m)\!+\!\chi(K)\!-\!\mu$.
II.3 Momentum distribution function
Once $G(K)$ is known, the momentum distribution function $n({\bm{k}})\ (=\langle c_{{\bm{k}}\sigma}^{+}c_{{\bm{k}}\sigma}\rangle)$ is calculated as
$n({\bm{k}})=T\sum_{\omega_{n}}G(K)e^{-i\omega_{n}0^{+}},$ (10)
where the Matsubara sum is taken by the procedure explained in Appendix A. Accuracy of $n({\bm{k}})$ may be checked by the three sum rules related to total electron number, total kinetic energy, and total kinetic-energy fluctuation, as derived in Ref. [54]. Those sum rules are conveniently expressed in terms of the $n$th-power integral $I_{n}$, given as
[…]
with $x\!=\!|{\bm{k}}|/k_{\rm F}$ and $n(x)\!=\!n({\bm{k}})$. The rigorous values for $I_{n}$ with $n\!=\!2$, $4$, and $6$ are: $I_{2}\!=\!1/3$, $I_{4}\!=\!1/5+\alpha^{2}r_{s}^{2}(-\varepsilon_{c}\!-\!r_{s}\varepsilon^{\prime})/3$, and $I_{6}\!=\!8/105\!+\!5I_{4}^{2}/3\!+\!5\alpha r_{s}[B(r_{s})\!-\!2/3\!+\!2g(0)/3]/(9\pi)$ with $g(0)$ the on-top value of the pair distribution function and $B(r_{s})$ specified in Eq. (51) in Ref. [54]. After several tentative calculations, we find that $n({\bm{k}})$ as obtained from Eq. (10) satisfies the $I_{2}$ sum rule up to 5 digits or more and $I_{4}$ up to 3 digits, but $I_{6}$ up to only a single digit in most cases, indicating that $n({\bm{k}})$ for $|{\bm{k}}|\!\gtrsim\!2k_{\rm F}$ is not accurate enough, which probably reflects the fact that the $k$-mesh (or grid) in ${\bm{k}}$ space in the iterative calculation of $G(K)$ is not dense enough for $|{\bm{k}}|\!\gtrsim\!2k_{\rm F}$.
As a remedy for this problem in $n({\bm{k}})$, we have modified $n({\bm{k}})\ [=\!n(x)]$ to $n_{c}(x)$ by the procedure explained in Appendix B in which the behavior in the region of $x\!\gtrsim\!2$ is rectified by the introduction of $n_{\rm IGZ}(x)$ obtained by the parametrization scheme described in Ref. [54]. In actual calculations, $n_{c}(x)$ always satisfies all the three sum rules up to at least 5 digits (and mostly 7-8 digits). As an example, the results of $n({\bm{k}})$, $n_{c}(x)$, and $n_{\rm IGZ}(x)$ are shown for $r_{s}=3.93$ in Fig. 10(a) in which we can barely see the difference among those three functions on the scale of the figure.
II.4 Polarization function
The charge response function $Q_{c}(Q)$ is related to the polarization function in the charge channel $\Pi(Q)$ through
$Q_{c}(Q)=-\frac{\Pi(Q)}{1+V({\bm{q}})\Pi(Q)},$ (12)
and the formal definition of $\Pi(Q)$ is written as
$\Pi(Q)$ $=-2\sum_{K}G(K)G(K\!+\!Q)\Gamma(K,K\!+\!Q)$
$\equiv-2\,T\sum_{\omega_{n}}\sum_{{\bm{k}}}G(K)G(K\!+\!Q)\Gamma(K,K\!+\!Q),$ (13)
where $\Gamma(K,K\!+\!Q)$ is the three-point vertex function in the charge channel.
In the many-body problem, it is often the case that $\Pi(Q)$ is less singular than $G(K)$. In fact, in 1D Luttinger liquids, $\Pi(Q)$ is easily obtained and does not exhibit any NFL-related singularity [12]. In 3D Fermi liquids, it is shown that no nonanalytic corrections are contained in $\Pi(Q)$, in sharp contrast to the polarization function in the spin channel in which a nonanalytic ${\bm{q}}^{2}\ln|{\bm{q}}|$-term exists [52]. Thus we can expect that in the 3D electron gas, even if there were a branch-cut singularity in $G(K)$ at low $T$, it would not induce any singular effects on $\Pi(Q)$, implying that $\Pi(Q)$ will be reliably determined even at $T$ of the order of $0.1\varepsilon_{\rm F}$. On this assumption, we consider $Q_{c}({\bm{q}},0)$ obtained by Monte Carlo simulations [110, 111, 83, 112, 113, 85, 114] as sufficiently accurate data, based on which various parametrized forms for the conventional charge local-field factor $G_{+}(Q)$ have been proposed [115, 116, 117, 118, 119, 120, 121].
In view of this situation, we regard $\Pi(Q)$ as a quantity precisely known from the outset in the self-consistent iteration loop. In actual calculations, $\Pi(Q)$ is given either with $G_{+}(Q)$ as
$\Pi(Q)=\frac{\Pi_{0}(Q)}{1-V({\bm{q}})G_{+}(Q)\Pi_{0}(Q)},$ (14)
or with $G_{s}(Q)$ due to Richardson and Ashcroft [115] as
$\Pi(Q)=\frac{\Pi_{\rm WI}(Q)}{1-V({\bm{q}})G_{s}(Q)\Pi_{\rm WI}(Q)},$ (15)
where the Lindhard function $\Pi_{0}(Q)$ is calculated as [122]
$\Pi_{0}(Q)$ $=-2\sum_{K}G_{0}(K)G_{0}(K\!+\!Q)$
$=4\int\frac{d^{3}k}{(2\pi)^{3}}\ n_{0}({\bm{k}})\frac{\varepsilon_{{\bm{k}}+{\bm{q}}}\!-\!\varepsilon_{\bm{k}}}{\omega_{q}^{2}\!+\!(\varepsilon_{{\bm{k}}+{\bm{q}}}\!-\!\varepsilon_{\bm{k}})^{2}},$ (16)
with $n_{0}({\bm{k}})\!=\!\theta(k_{\rm F}\!-\!|{\bm{k}}|)$ the step function and $\Pi_{\rm WI}(Q)$ is given by
$\Pi_{\rm WI}(Q)=4\int\frac{d^{3}k}{(2\pi)^{3}}\ n({\bm{k}})\frac{\varepsilon_{{\bm{k}}+{\bm{q}}}\!-\!\varepsilon_{\bm{k}}}{\omega_{q}^{2}\!+\!(\varepsilon_{{\bm{k}}+{\bm{q}}}\!-\!\varepsilon_{\bm{k}})^{2}}.$ (17)
Note that $G_{s}(Q)$ is obtained from $G_{+}(Q)$ as
$G_{s}(Q)=G_{+}(Q)+\frac{1}{V({\bm{q}})\Pi_{0}(Q)}\frac{\Pi_{0}(Q)\!-\!\Pi_{\rm WI}(Q)}{\Pi_{\rm WI}(Q)},$ (18)
but in this paper, we shall adopt the function form (a slightly modified Richardson-Ashcroft parametrization) prescribed in Eq. (58) in Ref. [54] for $G_{s}(Q)$.
In Eq. (17), we employ $n_{c}(x)$ for $n({\bm{k}})$ in order to make $\Pi_{\rm WI}({\bm{q}},0)$ correctly behave in the limit of $|{\bm{q}}|\to\infty$. On the other hand, the behaviors of $\Pi_{0}(Q)$ and $\Pi_{\rm WI}(Q)$ in the limit of $Q\to Q_{0}\!\equiv\!\{{\bm{0}},0\}$ can be derived directly from Eqs. (II.4) and (17), respectively; in the $\omega$-limit (i.e., ${\bm{q}}\to{\bm{0}}$ first, and then $\omega_{q}\to 0$), we obtain
$\Pi_{0}(Q)=\Pi_{\rm WI}(Q)=\frac{D_{\rm F}}{3}\frac{v_{\rm F}^{2}{\bm{q}}^{2}}{\omega_{q}^{2}},$ (19)
with $v_{\rm F}=k_{\rm F}/m$ and in the $q$-limit (i.e., $\omega_{q}\to 0$ first, and then ${\bm{q}}\to{\bm{0}}$), we obtain
$\Pi_{0}(Q)=D_{\rm F},\ \ {\rm and}\ \ \Pi_{\rm WI}(Q)=D_{\rm F}I_{0},$ (20)
where $I_{0}$ is defined in Eq. (11) with $n\!=\!0$. Combining the above-mentioned behavior of $\Pi_{0}(Q)$ with the constraints imposed on $G_{+}(Q)$ (or that of $\Pi_{\rm WI}(Q)$ with those on $G_{s}(Q)$) at $Q\to Q_{0}$, we see that in the $\omega$-limit,
$\Pi(Q)=\frac{D_{\rm F}}{3}\frac{v_{\rm F}^{2}{\bm{q}}^{2}}{\omega_{q}^{2}}=\frac{n{\bm{q}}^{2}}{m\omega_{q}^{2}},$ (21)
and in the $q$-limit,
$\Pi(Q)=D_{\rm F}\frac{\kappa}{\kappa_{\rm F}}=\frac{d\,n}{d\,\varepsilon_{\rm F}}\frac{d\,\varepsilon_{\rm F}}{d\,\mu}=\frac{d\,n}{d\,\mu}.$ (22)
The relations in Eqs. (21) and (22) are, respectively, known as the f-sum rule and the compressibility sum rule.
II.5 Vertex function
Formally, we can calculate $\Sigma(K)$ rigorously by
$\Sigma(K)\!=\!-\sum_{Q}W(Q)G(K\!+\!Q)\Gamma(K,K\!+\!Q),$ (23)
where $W(Q)$ is the effective interaction, given by
$W(Q)\!=\!\frac{V({\bm{q}})}{1\!+\!V({\bm{q}})\Pi(Q)}.$ (24)
Since we regard $\Pi(Q)$ as a precisely known quantity, $W(Q)$ is already known. As for $\Gamma(K,K\!+\!Q)$, we adopt the improved $GW\Gamma$ scheme described in Ref. [79]. According to Eqs. (54)-(58) in Ref. [79], $\Gamma(K,K\!+\!Q)$ is given in the product of two components as
$\Gamma(K,K\!+\!Q)=\widetilde{\Gamma}_{\rm WI}(K,K\!+\!Q)\Gamma_{\Pi}(K,K\!+\!Q),$ (25)
with
$\widetilde{\Gamma}_{\rm WI}(K,K\!+\!Q)\!=\!\frac{G^{-1}(K\!+\!Q)\!-\!G^{-1}(K)}{i\omega_{q}\!-\!(\varepsilon_{{\bm{k}}\!+\!{\bm{q}}}\!-\!\varepsilon_{{\bm{k}}}){\tilde{\eta}}_{1}(K\!+\!Q/2)},$ (26a)
$\Gamma_{\Pi}(K,K\!+\!Q)\!=\!\frac{\Pi(Q)}{\widetilde{\Pi}_{\rm WI}(Q)}$
$\hskip 59.75095pt+\frac{\Pi(Q)}{D_{\rm F}}\Biggl{\{}\frac{3i\omega_{q}}{v_{\rm F}^{2}{\bm{q}}^{2}}\biggl{[}-i\omega_{q}\left(1\!-\!\frac{\Pi_{1}(Q)}{\widetilde{\Pi}_{\rm WI}(Q)}\right)$
$\hskip 71.13188pt-\left(\varepsilon_{{\bm{k}}+{\bm{q}}}-\varepsilon_{\bm{k}}\right)\Bigl{(}1-{\tilde{\eta}}_{1}(K\!+\!Q/2)\Bigr{)}\biggr{]}$
$\hskip 71.13188pt+{\tilde{\eta}}_{2}(K;Q)-\frac{\Pi_{2}(Q)}{\widetilde{\Pi}_{\rm WI}(Q)}\Biggr{\}},$ (26b)
where the functional ${\tilde{\eta}}_{1}(K)$ is introduced as
${\tilde{\eta}}_{1}(K)$ $\equiv-\frac{\partial G^{-1}(K)}{\partial\varepsilon_{\bm{k}}}\bigg{/}\frac{\partial G^{-1}(K)}{\partial(i\omega_{n})},$ (27)
and three “modified polarization functionals”, $\widetilde{\Pi}_{\rm WI}(Q)$, $\Pi_{1}(Q)$, and $\Pi_{2}(Q)$, are, respectively, defined as
$\widetilde{\Pi}_{\rm WI}(Q)$ $\!\equiv\!2\!\sum_{K}\frac{G(K\!+\!Q)\!-\!G(K)}{i\omega_{q}\!-\!(\varepsilon_{{\bm{k}}\!+\!{\bm{q}}}\!-\!\varepsilon_{{\bm{k}}}){\tilde{\eta}}_{1}(K\!+\!Q/2)},$ (28a)
$\Pi_{1}(Q)$ $\!\equiv\!2\!\sum_{K}\frac{[G(K\!+\!Q)\!-\!G(K)](\varepsilon_{{\bm{k}}\!+\!{\bm{q}}}\!-\!\varepsilon_{{\bm{k}}})}{[i\omega_{q}\!-\!(\varepsilon_{{\bm{k}}\!+\!{\bm{q}}}\!-\!\varepsilon_{{\bm{k}}}){\tilde{\eta}}_{1}(K\!+\!Q/2)]i\omega_{q}},$ (28b)
$\Pi_{2}(Q)$ $\!\equiv\!2\!\sum_{K}\frac{[G(K\!+\!Q)\!-\!G(K)]{\tilde{\eta}}_{2}(K;Q)}{i\omega_{q}\!-\!(\varepsilon_{{\bm{k}}\!+\!{\bm{q}}}\!-\!\varepsilon_{{\bm{k}}}){\tilde{\eta}}_{1}(K\!+\!Q/2)}.$ (28c)
The functional ${\tilde{\eta}}_{2}(K;Q)$ will be specified later.
In Ref. [79], this functional form for $\Gamma(K,K\!+\!Q)$ was derived from the perspective of FLT, i.e., under the assumption of smooth $\Sigma(K)$ near the Fermi point $K_{\rm F}\!\equiv\!\{{\bm{k}}_{\rm F},0\}$ and consequently $G(K)$ in the form of Eq. (1). The final result in Eq. (25), together with Eqs. (26)-(28), however, does not explicitly contain any FLT-specific parameters such as the Landau interaction and $m^{*}$, implying that this functional form itself can be applied to NFL as well. In fact, ${\tilde{\eta}}_{1}(K)$, a key quantity leading to $m/m^{*}$ at $K\to K_{\rm F}$ in FLT, appears in this formalism as a direct consequence of the total current (or momentum) conservation law that should be satisfied even in NFL, indicating that ${\tilde{\eta}}_{1}(K)$ must also be a key quantity in NFL.
It is the most important advantage in this formalism that the Ward identity (WI) [81] is automatically satisfied, whatever choice one may make for ${\tilde{\eta}}_{1}(K)$ and ${\tilde{\eta}}_{2}(K;Q)$; if we choose ${\tilde{\eta}}_{1}(K)\!=\!{\tilde{\eta}}_{2}(K;Q)\!=\!1$, the present vertex function is nothing but the one in the original $GW\Gamma$ scheme [80] and $\widetilde{\Gamma}_{\rm WI}(K,K\!+\!Q)$ is reduced to the canonical form appearing in connection to WI as
$\Gamma_{\rm WI}(K,K\!+\!Q)\!=\!\frac{G^{-1}(K\!+\!Q)\!-\!G^{-1}(K)}{i\omega_{q}\!-\!\varepsilon_{{\bm{k}}\!+\!{\bm{q}}}\!+\!\varepsilon_{{\bm{k}}}}.$ (29)
Thus, we note that by just changing $\Gamma_{\rm WI}(K,K\!+\!Q)$ into $\widetilde{\Gamma}_{\rm WI}(K,K\!+\!Q)$, we can enter into a more advanced stage in which both WI and the total momentum conservation law are simultaneously fulfilled.
In actual numerical calculations, however, it turns out that it is not easy to proceed at each iteration step with ${\tilde{\eta}}_{1}(K)$ evaluated in accordance with Eq. (27). Because ${\tilde{\eta}}_{1}(K)$ reflects the total momentum conservation law only in its value at $K\!=\!K_{\rm F}$, we can approximate ${\tilde{\eta}}_{1}(K\!+\!Q/2)$ as $\eta_{1}(Q)$, a function depending only on $Q$ with the condition that $\eta_{1}(Q_{0})\!=\!{\tilde{\eta}}_{1}(K_{\rm F})$ in the $q$-limit. Under this approximation, we can reduce $\Gamma_{\Pi}(K,K\!+\!Q)$ into
$\Gamma_{\Pi}(K,K\!+\!Q)\!=\!\frac{\Pi(Q)}{\widetilde{\Pi}_{\rm WI}(Q)}\!+\!\frac{\Pi(Q)}{D_{\rm F}}\biggl{\{}\frac{1\!-\!\eta_{1}(Q)}{\eta_{1}(Q)}\frac{3i\omega_{q}}{v_{\rm F}^{2}{\bm{q}}^{2}}$
×$\hskip 1.42271pt\times\!\Bigl{[}i\omega_{q}\!-\!(\varepsilon_{{\bm{k}}\!+\!{\bm{q}}}\!-\!\varepsilon_{\bm{k}})\eta_{1}(Q)\Bigr{]}\!+\!{\tilde{\eta}}_{2}(K;Q)\!-\!\frac{\Pi_{2}(Q)}{\widetilde{\Pi}_{\rm WI}(Q)}\biggr{\}},$ (30)
and $\widetilde{\Pi}_{\rm WI}(Q)$ is easily calculated by
$\widetilde{\Pi}_{\rm WI}(Q)\!\equiv\!\widetilde{\Pi}_{\rm WI}({\bm{q}},i\omega_{q})=\frac{1}{\eta_{1}(Q)}\Pi_{\rm WI}\left({\bm{q}},\frac{i\omega_{q}}{\eta_{1}(Q)}\right),$ (31)
with $\Pi_{\rm WI}({\bm{q}},i\omega_{q})\!\equiv\!\Pi_{\rm WI}(Q)$ in Eq. (17).
Since we may take ${\tilde{\eta}}_{2}(K;Q)$ at our disposal, we choose it so as to make numerical calculations as easy as possible. By carefully inspecting the structure of each term in Eq. (30), we can think of a possible form for ${\tilde{\eta}}_{2}(K;Q)$ as
${\tilde{\eta}}_{2}(K;Q)=\eta_{1}(Q)\zeta_{1}(Q)\!+\!\Bigl{[}i\omega_{q}\!-\!(\varepsilon_{{\bm{k}}\!+\!{\bm{q}}}\!-\!\varepsilon_{\bm{k}})\eta_{1}(Q)\Bigr{]}$
×$\hskip 31.2982pt\times\frac{3}{v_{\rm F}^{2}{\bm{q}}^{2}}\Bigl{[}i\omega_{q}\zeta_{2}(Q)-(\varepsilon_{{\bm{k}}\!+\!{\bm{q}}}\!-\!\varepsilon_{\bm{k}})\zeta_{3}(Q)\Bigr{]},$ (32)
with arbitrary functions $\zeta_{i}(Q)$ ($i\!=\!1,2$, and $3$). By substituting this ${\tilde{\eta}}_{2}(K;Q)$ into Eq. (30), we immediately find the following: (i) $\zeta_{1}(Q)$ is irrelevant, because it is always cancelled by the corresponding term in $\Pi_{2}(Q)/\widetilde{\Pi}_{\rm WI}(Q)$. (ii) By choosing $\zeta_{2}(Q)$ as $[\eta_{1}(Q)-1]/\eta_{1}(Q)$, we can eliminate the first term in the curly brackets in Eq. (30). (iii) Because the term containing $\zeta_{3}(Q)$ is rather difficult to treat, it would be better to approximate $(\varepsilon_{{\bm{k}}\!+\!{\bm{q}}}\!-\!\varepsilon_{\bm{k}})\zeta_{3}(Q)$ by $v_{\rm F}q\zeta_{3}(Q)$ with $q=|{\bm{q}}|$. As a result, we have arrived at the functional form for $\Gamma(K,K\!+\!Q)$, written as
$\Gamma(K,K\!+\!Q)\!=\!\frac{G^{-1}(K\!+\!Q)\!-\!G^{-1}(K)}{i\omega_{q}\!-\!(\varepsilon_{{\bm{k}}\!+\!{\bm{q}}}\!-\!\varepsilon_{\bm{k}})\eta_{1}(Q)}\,\frac{\Pi(Q)}{\widetilde{\Pi}_{\rm WI}(Q)}$
$\hskip 28.45274pt\!-\!\bigl{[}G^{-1}(K\!+\!Q)\!-\!G^{-1}(K)\bigr{]}\frac{\Pi(Q)}{D_{\rm F}}\frac{3\zeta_{3}(Q)}{v_{\rm F}q}.$ (33)
It should be noted that we can rigorously reproduce $\Pi(Q)$ irrespective of $\eta_{1}(Q)$ and $\zeta_{3}(Q)$ in substituting this functional form for $\Gamma(K,K\!+\!Q)$ into Eq. (II.4), verifying the internal consistency of our formulation.
In order to determine $\eta_{1}(Q)$ and $\zeta_{3}(Q)$, we need to consider several constraints to be imposed on them so as to make $\Gamma(K,K\!+\!Q)$ behave correctly in various limits. In line with such constraints as discussed in Appendix C, we will use $\eta_{1}(Q)$ in Eq. (C3) and $\zeta_{3}(Q)$ given by
$\zeta_{3}(Q)\!=\!-\frac{1}{3}\,\frac{\Pi_{0}(Q)\!-\!\Pi_{\rm WI}(Q)[1\!-\!\beta_{3}G_{s}(Q)]}{\Pi_{\rm WI}(Q)},$ (34)
in this paper.
Figure 2: In (a), $\eta_{1}({\bm{q}},0)$ at $T\!=\!10^{-4}\varepsilon_{\rm F}$ is plotted in the range of $0\!<\!q\!<\!10k_{\rm F}$ for the 3D electron gas with the density region corresponding to simple metals and in (b), the same function is drawn with $q$ in an enlarged scale near $q=0$.
In Fig. 2, the self-consistently determined results of $\eta_{1}({\bm{q}},0)$ are plotted at $T\!=\!10^{-4}\varepsilon_{\rm F}$ for the 3D homogeneous electron gas with $62\!<\!r_{s}\!<\!$ 6. As seen from this figure, $\eta_{1}(Q)$ is smoothly converged to unity for $q\gtrsim 10k_{\rm F}$, but it exhibits a rather rapid change near $q\approx 0$, indicating that $\Sigma(K)$ is not smooth enough for $K\approx K_{\rm F}$, a symptom of the possible breakdown of FLT. Incidentally, in obtaining $\eta_{1}(Q)$, we need to know the static physical quantities, $E({\bm{k}},0)$ and $Z({\bm{k}},0)$, which can be obtained by an extrapolation procedure explained in Appendix D.
In Eq. (34), the term involving the parameter $\beta_{3}$ is introduced so as to rigorously reproduce $\mu_{c}$, the value accurately given in Eq. (4). The appropriately determined values for $\beta_{3}$ providing the correct $\mu_{c}$ with $r_{s}$ in the region of $61<r_{s}<$ 6 are shown in Fig. 3 in which we take $T$ as $10^{-4}\varepsilon_{\rm F}$. As seen from the figure, the magnitude of $\beta_{3}$ is small, i.e., of the order of 0.02, letting us know that the $\beta_{3}$ term exerts only limited effects on $\Sigma(K)$.
Figure 3: Results of $\beta_{3}$ in Eq. (34) determined so as to reproduce the accurate $\mu_{c}$ for the 3D electron gas with the density parameter $r_{s}$ in the range of $61\!<\!r_{s}\!<\!$ 6.
II.6 Self-energy
By substituting Eq. (33) with Eq. (34) into Eq. (23), we obtain an expression for the calculation of $\Sigma(K)$ as
$\Sigma(K)=\Sigma_{a}+\Sigma_{b}(K)+\Lambda(K)G^{-1}(K),$ (35)
with
$\Sigma_{a}\ =-\sum_{Q}W_{\rm WI}(Q)R_{\rm WI}(Q)$
$\hskip 17.07182pt=-T\sum_{\omega_{q}}\int_{0}^{\infty}\frac{q^{2}dq}{2\pi^{2}}\,W_{\rm WI}(Q)R_{\rm WI}(Q),$ (36a)
$\Sigma_{b}(K)\!=\!-\sum_{K^{\prime}}\overline{W}_{\rm WI}(Q)G(K^{\prime})\overline{\Gamma}_{\rm WI}(K,K^{\prime})$
$\hskip 28.45274pt=-T\sum_{\omega_{n^{\prime}}}\frac{1}{k}\int_{0}^{\infty}\frac{qdq}{4\pi^{2}}\,\overline{W}_{\rm WI}({\bm{q}},i\omega_{n^{\prime}}\!-\!i\omega_{n})$
×$\hskip 51.21504pt\times\int_{|k-q|}^{k+q}k^{\prime}dk^{\prime}G(K^{\prime})\overline{\Gamma}_{\rm WI}(K,K^{\prime}),$ (36b)
$\Lambda(K)\,=\!\sum_{K^{\prime}}W_{\rm WI}(Q)R_{\rm WI}(Q)G(K^{\prime})$
$\hskip 28.45274pt=T\sum_{\omega_{n^{\prime}}}\frac{1}{k}\int_{0}^{\infty}\frac{qdq}{4\pi^{2}}\,W_{\rm WI}({\bm{q}},i\omega_{n^{\prime}}\!-\!i\omega_{n})$
[…]
where $W_{\rm WI}(Q)$, $\overline{W}_{\rm WI}(Q)$, $R_{\rm WI}(Q)$, and $\overline{\Gamma}_{\rm WI}(K^{\prime},K)$ are, respectively, defined as
$W_{\rm WI}(Q)=W(Q)\frac{\Pi(Q)}{\Pi_{\rm WI}(Q)}$
$\hskip 36.98866pt=\frac{V({\bm{q}})}{1+V({\bm{q}})\Pi_{\rm WI}(Q)[1-G_{s}(Q)]},$ (37a)
$\overline{W}_{\rm WI}(Q)=W_{\rm WI}(Q)\,\frac{\Pi_{\rm WI}(Q)}{\widetilde{\Pi}_{\rm WI}(Q)},$ (37b)
$R_{\rm WI}(Q)\ =\,\frac{\Pi_{0}(Q)\!-\!\Pi_{\rm WI}(Q)[1-\beta_{3}G_{s}(Q)]}{D_{\rm F}v_{\rm F}q},$ (37c)
$\overline{\Gamma}_{\rm WI}(K,K^{\prime})=\frac{G^{-1}(K^{\prime})\!-\!G^{-1}(K)}{i\omega_{q}\!-\!(\varepsilon_{\bm{k}^{\prime}}\!-\!\varepsilon_{\bm{k}})\eta_{1}(Q)}.$ (37d)
Here, $\Sigma_{a}$ is independent of $K$ and directly connected to the chemical potential shift, $\Sigma_{b}(K)$ is the main contribution to the self-energy, and $\Lambda(K)$ partially contributes to the renormalization factor.
II.7 Self-consistent iteration loop
To summarize this section, Fig. 4 schematically displays a self-consistent iteration loop to determine $\Sigma(K)$ on the imaginary $\omega_{n}$-axis. In the actual numerical implementation, instead of $\Sigma(K)$, we calculate a couple of real functions in Eq. (9), $Z(K)$ and $E(K)$, both of which are even in $\omega_{n}$ and depend on ${\bm{k}}$ only through $k\ (\equiv|{\bm{k}}|)$. Those functions will be evaluated only at a finite number of points in $(k,\omega_{n})$-space, $\{k_{i},\omega_{j}\}$, and if we need the values of those functions at $K$ other than those selected points, we will employ the two-dimensional cubic spline interpolation method. In Appendix E, we make a more detailed explanation of $\{k_{i},\omega_{j}\}$, together with some remarks on the numerical integration in Eqs (36a)-(36c) and the numerical analytic continuation.
Figure 4: Self-consistent iteration loop to determine the self-energy $\Sigma(K)$ in the present calculation scheme.
The actual implementation of this iteration loop starts with $\Sigma_{0}(K)$ in RPA, given by
$\Sigma_{0}(K)=-\sum_{Q}\frac{V({\bm{q}})}{1+V({\bm{q}})\Pi_{0}(Q)}\,G_{0}(K+Q)$ (38)
and ends up when the relative difference in $\Sigma(K)$ between input and output at each selected point becomes less than $10^{-5}$. In revising the input $\Sigma(K)$ at each step during the iteration loop, we employ the second Broyden’s method [123, 124, 125]. We need $10-100$ iteration steps depending on $r_{s}$ and $T$ to obtain converged results for $\Sigma(K)$. The calculated $\Sigma(K)$ is converted into the retarded self-energy $\Sigma^{R}({\bm{k}},\omega\!+\!i\gamma)$ with $\gamma\!=\!\pi T$ through numerical analytic continuation with the use of Padé approximants [126].
III Calculated Results
III.1 One-Particle spectral function
[…]
In Fig. 5, the one-particle spectral function $A({\bm{k}},\omega)\ [\equiv\!-{\rm Im}G^{R}({\bm{k}},\omega\!+\!i\gamma)/\pi]$ is plotted as a function of $\omega$ for $r_{s}\!=\!3.93$, corresponding to sodium, at $T\!=\!4\!\times\!10^{-3}\varepsilon_{\rm F}$ with $k\ (\equiv\!|{\bm{k}}|)\!=\!0.0-2.2$ in units of $k_{\rm F}$. This shows a well-known typical behavior, characterized by the dominant quasiparticle peak and the associated one-plasmon satellites. In fact, this result is essentially the same, even quantitatively, as the one given at $r_{s}=4$ in Fig. 3(a) in Ref. [79] in which both ${\tilde{\eta}}_{1}$ and ${\tilde{\eta}}_{2}$ were taken as unity. Therefore, we cannot find any indication of the breakdown of FLT at least for sodium at $T=4\!\times\!10^{-3}\varepsilon_{\rm F}$, as is the case in all previous studies on alkali metals.
By decreasing $T$ down to $10^{-4}\varepsilon_{\rm F}$, however, we find a totally new situation in the low-energy region of $\omega$ for $k\sim k_{\rm F}$ at $r_{s}=3.93$, as shown in Fig. 6. Although nothing special is seen at $k\!=\!k_{\rm F}$ at all temperatures, a shoulder or bump structure begins to develop at $T\!=\!10^{-3}\varepsilon_{\rm F}$ and a clear peak structure emerges at $T\!=\!10^{-4}\varepsilon_{\rm F}$ for $k$ not equal to $k_{\rm F}$ but close to it. The energy of this new peak is about a half of that of the corresponding quasiparticle peak at the same $k$ and thus a new mode, if any, associated with this peak (to be dubbed “excitron”) is characterized by the very low excitation energy of the order of $0.1\varepsilon_{\rm F}$ or less.
Figure 6: Change of $A({\bm{k}},\omega)$ with the decrease of $T$ from $4\!\times\!10^{-3}$ down to $10^{-4}$ in units of $\varepsilon_{\rm F}$ for $r_{s}=3.93$ with $k=0.92k_{\rm F}$ (in blue), $k_{\rm F}$ (in gray), and $1.08k_{\rm F}$ (in red).
In Fig. 7, we show the change in shape and position of this new peak with the increase of $k$ from $0.95k_{\rm F}$ to $1.05k_{\rm F}$ through $k_{\rm F}$ for $r_{s}\!=\!3.93$ at $T\!=\!10^{-4}\varepsilon_{\rm F}$, from which we see that this new peak is absorbed into (or perfectly overlapped with) the dominant quasiparticle peak at $k=k_{\rm F}$. Thus, we need to keep $k$ away from $k_{\rm F}$ to detect this new peak, but at the same time, as $|k-k_{\rm F}|$ increases, the peak height decreases, making the detection rather difficult. To compromise between these competing factors, it would be good to search for this peak at $k$ in the range $|k-k_{\rm F}|\sim 0.03k_{\rm F}-0.2k_{\rm F}$. We have also calculated $A({\bm{k}},\omega)$ at $T\!=\!10^{-4}\varepsilon_{\rm F}$ for other values of $r_{s}$ in the range of $61<r_{s}<$ 6 to find that this new peak always appears with qualitatively the same features as those described above, including the behavior with the change of $k$ and $T$, though quantitatively the peak strength (or height) depends rather strongly on $r_{s}$; as $r_{s}$ increases, the peak emerges more vividly and strongly. A more detailed analysis on the character of this new peak (or excitron) will be made in Sec. IV.
Figure 7: Change in shape and position of the new peak in $A({\bm{k}},\omega)$ with the increase of $k$ from $0.95k_{\rm F}$ to $1.05k_{\rm F}$ for $r_{s}=3.93$ at $T=10^{-4}\varepsilon_{\rm F}$. The new peak position is indicated by an arrow at each $k$.
The dispersion relation of excitron (or the peak position of this new mode), $\xi_{\bm{k}}$, at $T\!=\!10^{-4}\varepsilon_{\rm F}$ for $r_{s}=3.93$ and $5.20$ corresponding to Na and Rb, respectively, is drawn in Fig. 8, together with the quasiparticle dispersion relation ${\tilde{\varepsilon}}_{\bm{k}}$ determined by the quasiparticle peak position and the bare dispersion $\varepsilon_{\rm k}$. Although the results for $r_{s}=4.86$ corresponding to K are not shown here, they enter just between those of Na and Rb. For $|k-k_{\rm F}|\gtrsim 0.25k_{\rm F}$, $\xi_{\bm{k}}$ is not shown, because the new peak in either $k\lesssim 0.75k_{\rm F}$ or $k\gtrsim 1.25k_{\rm F}$ becomes broad and its peak height is very low, makig us very difficult to identify the peak position or the peak itself. For $k\sim k_{\rm F}$, the excitron dispersion $\xi_{\bm{k}}$ is linear and can be written as $\xi_{\bm{k}}=v_{\rm excitron}(k-k_{\rm F})$, but for $k$ not in the vicinity of $k_{\rm F}$, $\xi_{\bm{k}}$ deviates from this linear relation. The ratio of $v_{\rm excitron}/v_{\rm F}$ is plotted as a function of $r_{s}$ in Fig. 12 in Sec. III.4.
Figure 8: Dispersion relation $\xi_{\bm{k}}$ of the new peak (excitron) in comparison with the quasiparticle dispersion relation ${\tilde{\varepsilon}}_{\bm{k}}$ and the bare dispersion $\varepsilon_{\rm k}$ (the dotted curve) for $r_{s}\!=\!3.93$ and $5.20$ corresponding to Na and Rb, respectively, at $T\!=\!10^{-4}\varepsilon_{\rm F}$. For comparison, ${\tilde{\varepsilon}}_{\bm{k}}$ at $r_{s}\!=\!4$ in diffusion Monte Carlo (DMC) simulations is given by the brown solid curve.
The quasiparticle effective mass $m^{*}$ estimated by ${\tilde{\varepsilon}}_{\bm{k}}$ with $k$ in the range of $|k-k_{\rm F}|\lesssim 0.1k_{\rm F}$ is about the same as $m$, but it becomes smaller than $m$ for $k$ outside of this range. This change of $m^{*}$ with $k$ can easily be understood by the fact that $m^{*}$ is determined by the competition of exchange and correlation effects; the former makes $m^{*}$ small as understood by the fact that $m^{*}\!\to\!0$ at $k\!=\!k_{\rm F}$ in the exchange-only (or Hartree-Fock) approximation, while the latter makes $m^{*}$ large as easily guessed by just considering the heavy-fermion physics. Because the exchange effect is evaluated in first-order perturbation theory (and thus without energy denominators), it persists even for $k$ far away from $k_{\rm F}$. This is not the case for the correlation effect to which the second- and higher-order perturbation terms contribute. Thus, as $k$ goes away from the Fermi level, the correlation effect becomes weaker than that of exchange, making $m^{*}$ smaller than $m$. In this way, the dispersion relation ${\tilde{\varepsilon}}_{\bm{k}}$ can never be parabolic in the whole region of $k$ from 00 to $k_{\rm F}$, making the occupied bandwidth wider than that of the free-electron band. This widened bandwidth is also seen in the diffusion Monte Carlo (DMC) simulations [127, 94] as shown by the brown curve in Fig. 8; the magnitude of bandwidth in DMC is about the same as that in the present calculation, though its behavior of ${\tilde{\varepsilon}}_{\bm{k}}$ for $k\!\sim\!k_{\rm F}$ is much different, providing $m^{*}$ much smaller than $m$ which is not correct as we shall see in Sec. III.4.
In contrast to this widening of the computed occupied bandwidth, much narrower bandwidths have been observed by ARPES experiments [90, 93]. To account for this discrepancy, various explanations were proposed; in particular, the effect of final states in the optical measurements attracted attention [128, 80], but very recently yet another plausible explanation was proposed by emphasizing the importance of local dynamical correlations associated with an atom based on a more realistic model beyond the electron gas [129].
Figure 9: Change of $A({\bm{k}},\omega)$ for $k$ in the range from $0.92k_{\rm F}$ to $1.08k_{\rm F}$ at $r_{s}\!=\!3.93$ and $T\!=\!10^{-4}\varepsilon_{\rm F}$ with paying special attention to one- and two-plasmon satellites, indicated by blue smaller and red larger arrows, respectively.
In Fig. 9, $A({\bm{k}},\omega)$ is plotted on the hundred times wider scale of $\omega$ for $k$ in the range of $0.92-1.08$ in units of $k_{\rm F}$ at $r_{s}\!=\!3.93$ and $T\!=\!10^{-4}\varepsilon_{\rm F}$. On this large (not logarithmic but linear) scale, the excitron peak is not seen well as a separate structure from the dominant quasiparticle peak even at this very low $T$, but we can detect the structures associated with the energies of the order of $\varepsilon_{\rm F}$ instead. In fact, we can easily find the one-plasmon satellites (shown by blue smaller arrows) and also even the two-plasmon ones (shown by red larger arrows) which are usually not seen at all in GW-type calculations. As far as the author knows, this is the second theoretical study on successfully producing this two-plasmon satellites [106] without resort to the cumulant-related approximations [100, 101, 102, 103, 104, 105]. More details on this important achievement may be published elsewhere in the future, along with data for wider ranges of $k$ and comparisons with other related works.
III.2 Momentum distribution function
The momentum distribution function $n({\bm{k}})$ is calculated for $r_{s}$ in the range of $2.07-5.62$ at $T=10^{-4}\varepsilon_{\rm F}$ in accordance with the prescription described in Sec. II.3, together with Appendix B, in which three functions, $n({\bm{k}})$ defined in Eq. (10), $n_{\rm IGZ}(k/k_{\rm F})$, and $n_{c}(k/k_{\rm F})$, are introduced. As shown in Fig. 10(a), we cannot see the difference among those three functions at $r_{s}=3.93$ on the scale of this figure. In fact, as long as $r_{s}\lesssim 5.0$, those three functions provide virtually the same result. Even for $r_{s}>5$, a small difference between $n({\bm{k}})$ and $n_{c}(k/k_{\rm F})$ appears only for $k\gtrsim 1.5k_{\rm F}$. Therefore, in Fig. 10(b), we give only the results of $n_{c}(k/k_{\rm F})$ which satisfies all the three sum rules associated with $I_{2}$, $I_{4}$, and $I_{6}$ very accurately up to seven digits. We have also calculated $n({\bm{k}})$ at several other temperatures up to $4\times 10^{-3}\varepsilon_{\rm F}$ to find that the obtained results are virtually independent of $T$. Incidentally, the present results of $n_{c}(k/k_{\rm F})$ are essentially the same as those of the momentum distribution function given in Sec. II F in Ref. [54].
Figure 10: Momentum distribution function $n({\bm{k}})$ at $T\!=\!10^{-4}\varepsilon_{\rm F}$ for (a) $r_{s}\!=\!3.93$ corresponding to sodium and (b) $r_{s}$ in the range of $2.07-5.62$. In (a), data at $r_{s}\!=\!3.99$ in QMC are also included for comparison.
For comparison, the results of $n({\bm{k}})$ for $r_{s}=3.99$ in quantum Monte Carlo (QMC) simulations [96, 130] are depicted in Fig. 10(a). The data in QMC are seen to be in reasonably good agreement with our present ones, but we do not regard the QMC data as sufficiently accurate ones for the following reasons: (i) The QMC data are not verified to satisfy the three sum rules. In fact, by just looking at the difference between $n({\bm{k}})$ in QMC and $n_{c}(k/k_{\rm F})$ perfectly satisfying the three sum rules, we would consider that the QMC data might satisfy the $I_{2}$ sum rule, but they never satisfy other sum rules. (ii) Although the data in Ref. [130] (green diamonds) was claimed to be the same as those in Ref. [96] (blue dotted-dashed curve), they are actually very different, as clearly seen in Fig. 10(a), indicating that the size extrapolation, an inevitable process in QMC to obtain physical quantities in the bulk system, is not reliable enough to produce a definite and well-converged result.
III.3 Quasiparticle Renormalization factor
In all preceding works in which FLT was assumed to be valid, the quasiparticle renormalization factor $z^{*}$ is nothing but $Z({\bm{k}}_{\rm F},0)^{-1}$. In the present study, however, the quasiparticle peak is overlapped with that of excitron at $k=k_{\rm F}$ as seen in Fig. 7, implying a possibility that $z^{*}$ differs from $Z({\bm{k}}_{\rm F},0)^{-1}$. Because we will come to know in Sec. IV that the singularity associated with the excitron is well described by a branch cut, we do not expect any contribution from the excitron to the jump of $n({\bm{k}})$ at the Fermi level, suggesting us to consider this jump as $z^{*}$. At the same time, we may regard the difference between $Z({\bm{k}}_{\rm F},0)^{-1}$ and the jump as the strength of the excitron peak at the Fermi level, $\delta z$, namely,
$z^{*}$ $\equiv n_{c}(1\!-\!0^{+})-n_{c}(1\!+\!0^{+}),$ (39)
$\delta z$ $\equiv\frac{1}{Z({\bm{k}}_{\rm F},0)}-z^{*}.$ (40)
In actual calculations, we find that $\delta z$ defined in Eq. (40) is always positive, which is consistent with our interpretation of $\delta z$ from a physical point of view.
Figure 11: Quasiparticle renormalization factor $z^{*}$ as a function of $r_{s}$ given by the green solid curve with circles at $T\!=\!10^{-4}\varepsilon_{\rm F}$. For comparison, the results in $G_{0}W_{0}$ (blue dashed curve) and in EPX (purple dotted-dashed curve) are shown, together with the data in QMC simulations (big brown squares) and the experimental data (solid circles with error bars) for Al, Li, and Na. The data for $\delta z$ representing the strength of the excitron peak (red solid curve with squares) are also given, together with those of $n_{0}\equiv n_{c}(0)$ and $n_{\pm}\equiv n_{c}(1\!\pm\!0^{+})$ by dotted curves.
In Fig. 11, we plot our results of $z^{*}$ and $\delta z$ as a function of $r_{s}$ by green solid curve with circles and red solid curve with squares, respectively. For reference, the data for $n_{0}\ [=\!n_{c}(0)]$ and $n_{\pm}\ [=\!n_{c}(1\!\pm\!0^{+})]$ are also given by the black dotted curves. This figure clearly shows that $z^{*}$ is larger than $\delta z$ by $50\!-\!100$ times, implying that the excitron will exert its effect, if any, on bulk physical quantities by only a very small amount for $r_{s}<6$.
For comparison, the preceding results of $z^{*}$ in both experiments and theories are added to the figure: Compton-scattering studies were done on Al [98], Li [131, 97], and Na [99] and the obtained results are indicated by the big black solid circles with error bars, while the data in QMC [96, 97] are by the big brown squares. For the sake of reference, the results of $z^{*}$ in $G_{0}W_{0}$ and EPX methods are, respectively, shown by the blue dashed and purple dotted-dashed curves. We find that (i) our results of $z^{*}$ perfectly reproduce those experimental ones, (ii) they are also in very good agreement with the QMC data, and (iii) they are virtually the same as the old data in EPX, confirming the accuracy of the results in Ref. [132].
In the literature, it is often the case that the results of $z^{*}$ are given to be much higher than our present results or even those in $G_{0}W_{0}$, but those results are not correct, simply because such results originate from an inappropriate treatment of the correlation effect near the Fermi level where the energy denominators diverge. As discussed in rather details in Ref. [132], any theoretical frameworks without correctly taming the divergent energy denominators will fail to produce correct values of $z^{*}$. To be more concrete, a theory with the use of Jastrow-type variational trial functions is, in general, not a good choice. Even in QMC or DMC, if those simulations start with Jastrow-based variational Monte Carlo, final results may inherit the demerits of Jastrow functions. This might be one of the reasons why DMC does not provide correct ${\tilde{\varepsilon}}_{\bm{k}}$ near the Fermi level in Fig. 8.
III.4 Quasiparticle Effective mass
The quasiparticle effective mass $m^{*}$ can be determined through the derivative of ${\tilde{\varepsilon}}_{\bm{k}}$ at $k\!=\!k_{\rm F}$ and the obtained result is drawn by the green solid curve with circles in Fig. 12. Our present result is very close to that of Simion and Giuliani [74] given by the black dashed curve and it is also in excellent agreement with the very recent data provided by diagrammatic Monte Carlo calculation [84] (blue diamonds) and QMC [86] (brown squares). The result in $G_{0}W_{0}$ shown by the black dotted curve is not quite the same as ours, but it may still be regarded as a semi-quantitatively good result. This success of $G_{0}W_{0}$ probably reflects the fact that ${\tilde{\varepsilon}}_{\bm{k}}$ itself is well reproduced by $G_{0}W_{0}$ due to the strong mutual cancellation between the self-energy effect and the vertex correction as a consequence of the Ward identity [87]. Thus, if we are not concerned with the physics of highly correlated phenomena such as excitron, $G_{0}W_{0}$ is a good choice for many purposes, because it is computationally very cheap yet provides qualitatively correct results.
Figure 12: Quasiparticle effective mass in units of the free-electron mass $m^{*}/m$ as a function of $r_{s}$, together with the excitron velocity in units of $v_{\rm F}$. The results in $G_{0}W_{0}$, the method by Simion-Giuliani, diagrammatic Monte Carlo (DiagMC), and QMC are also shown for comparison.
Compared with the quasiparticle velocity at the Fermi level $v^{*}_{\rm F}\!\equiv\!(m/m^{*})v_{\rm F}\!\approx\!v_{\rm F}$, the velocity of excitron at the Fermi level $v_{\rm excitron}$, which can be obtained through the derivative of $\xi_{\bm{k}}$ at $k\!=\!k_{\rm F}$, is found to be typically about a half of $v^{*}_{\rm F}$, as seen in Fig. 12 in which $v_{\rm excitron}$ is shown in units of $v_{\rm F}$ by the red solid curve with squares.
IV Details of Excitron
So far, the excitron is introduced only as a new low-energy peak in $A({\bm{k}},\omega)$, but here we try to understand its features in terms of the self-energy, either the retarded one $\Sigma^{R}({\bm{k}},\omega\!+\!i\gamma)$ or the thermal one $\Sigma({\bm{k}},i\omega_{n})$. For this purpose, we focus exclusively on the case of $r_{s}\!=\!3.93$ at $T\!=\!2\times\!10^{-4}\varepsilon_{\rm F}$ in Secs. IV.1--IV.4, partly because this is a typical example exhibiting a clear excitron peak in $A({\bm{k}},\omega)$ and partly because we can obtain a completely convergent result of $\Sigma(K)$ much more easily in this system than in those at larger $r_{s}$ and/or lower $T$. In Sec. IV.5, we examine the $T$-dependence of $\Sigma^{R}({\bm{k}},\omega\!+\!i\gamma)$ by changing $T$ in the range of $(1-8)\times 10^{-4}\varepsilon_{\rm F}$.
[…]
Figure 13: Peaks in the one-particle spectral function $A({\bm{k}},\omega)$ in (a) and the corresponding structure in the retarded self-energy $\Sigma^{R}({\bm{k}},\omega\!+\!i\gamma)$ in (b) at $k\!\equiv\!|{\bm{k}}|\!=\!0.95k_{\rm F}$ for the 3D homogeneous electron gas at $r_{s}\!=\!3.93$ and $T\!=\!2\times\!10^{-4}\varepsilon_{\rm F}$. The region of $\omega\sim 0$ is enlarged to show the behavior leading to the quasiparticle and excitron peaks in the inset in (b).
In Fig. 13(a), $A({\bm{k}},\omega)$ in this system is drawn in a logarithmic scale as a function of $\omega$ at $k\!\equiv\!|{\bm{k}}|\!=\!0.95k_{\rm F}$. For $\omega$ in the range $(-3\varepsilon_{\rm F},3\varepsilon_{\rm F})$, there exist four peaks in $A({\bm{k}},\omega)$ and the corresponding structure in $\Sigma^{R}({\bm{k}},\omega\!+\!i\gamma)\!-\!\mu_{x}\!-\!\mu_{c}$ is given in Fig. 13(b); at the quasiparticle peak position, both real and imaginary parts in $\Sigma^{R}({\bm{k}},\omega\!+\!i\gamma)\!-\!\mu_{x}\!-\!\mu_{c}$ vary very smoothly with vanishingly small magnitudes, as shown in the inset, in accordance with the assumption in FLT. On the other hand, the one-plasmon satellites are associated with large variations in both real and imaginary parts in the shape of functions as can be found in the Lorentz oscillator model, reflecting an electron motion in the electric field induced by the plasmon. At the excitron peak position, although the magnitudes of the variations are by far small, about a one-hundredth, the real and imaginary parts behave in a way very similar to those at the one-plasmon satellites, implying that the excitron peak must also be connected with the motion of an electron in the field induced by some kind of low-energy (of the order of $0.1\varepsilon_{\rm F}$) excitations.
Figure 14: Two-dimensional contour map of the dynamical structure factor $S(q,\omega)$ for the electron gas characterized by the same parameters as those in Fig. 13.
In order to pinpoint the relevant low-energy excitations to bring about the excitron, we show the calculated result of the structure factor $S({\bm{q}},\omega)$, defined by
$S({\bm{q}},\omega)=-\frac{1}{\pi}\frac{1}{1-e^{-\omega/T}}\,{\rm Im}\,Q_{c}^{R}({\bm{q}},\omega),$ (41)
in Fig. 14 to demonstrate that, though the plasmon contribution in the range of $\omega\approx 1.9-2.4\varepsilon_{\rm F}$ with $q\lesssim k_{\rm F}$ overwhelmingly dominates, the most important contribution in the range of $\omega\approx 0-0.2\varepsilon_{\rm F}$ comes from the tightly-bound electron-hole pair excitations with $q\!\approx\!2k_{\rm F}$, the region indicated by the yellow shaded area in the figure. Incidentally, the result for $S({\bm{q}},\omega)$ in Fig. 14 is virtually the same as that given for $r_{s}=4$ in Fig. 1(a) in Ref. [54], basically because the present result for $Q_{c}^{R}({\bm{q}},\omega)$ is essentially the same as those in the previous publications, not only in Ref. [54] but also in Refs. [133, 134].
Now, our task is to study the electron-electron interaction mediated by the tightly-bound electron-hole pair excitations $V_{\rm ex}(K,K^{\prime};Q)$ and its contribution to the self-energy $\Sigma_{\rm ex}(K)$; diagrammatically, they are given in Fig. 15(a) and Fig. 15(c), respectively, with the electron-hole irreducible (defined in the horizontal view) four-point interaction $\tilde{I}(K_{1},K_{2};K_{1}\!+\!Q,K_{2}\!-\!Q)$ in Fig. 15(b). For $Q\approx 2K_{\rm F}$ and an electron on the Fermi sphere, i.e., $|{\bm{k}}|\!\approx k_{\rm F}$, the dominant contribution to $\Sigma_{\rm ex}(K)$ comes from the scattered states $K\!+\!Q$ also on the Fermi sphere. Since $|{\bm{q}}|\!\approx\!2k_{\rm F}$, this is only possible for ${\bm{k}}\!+\!{\bm{q}}\!\approx\!-{\bm{k}}$ and ${\bm{q}}\approx-2{\bm{k}}$. A similar restriction also applies to each pair polarization process in $V_{\rm ex}(K,K+Q;Q)$ in Fig. 15(a); the important contribution arises only for ${\bm{k}}_{n}\!\approx\!{\bm{k}}$ in $G(K_{n})$ and ${\bm{k}}_{n}\!+\!{\bm{q}}\!\approx\!-{\bm{k}}$ in $G(K_{n}\!+\!Q)$. The above restrictions clearly indicate that once we choose an electron characterized by ${\bm{k}}$, all the processes of its scatterings and the associated electron-hole pair excitations occur predominantly in either parallel or antiparallel to the ${\bm{k}}$-direction.
Figure 15: Diagrammatic representation of $V_{\rm ex}(K,K^{\prime};Q)$ the electron-electron attractive interaction induced by multiple excitations of tightly-bound electron-hole pairs in (a), with the electron-hole irreducible four-point interaction $\tilde{I}$ in (b), and its contribution to the self-energy $\Sigma_{\rm ex}(K)$ in (c).
If we approximate $G(K_{n})$ and $G(K_{n}\!+\!Q)$ by $G_{0}(K_{n})$ and $G_{0}(K_{n}\!+\!Q)$, respectively, in Fig. 15(a), we can easily obtain an approximate expression for $V_{\rm ex}$ as
$V_{\rm ex}\approx-\frac{V({2\bm{k}}_{\rm F})^{2}\Pi_{0}(2K_{\rm F})}{1\!\!-\tilde{I}(K_{\rm F},-K_{\rm F};-K_{\rm F},K_{\rm F})\Pi_{0}(2K_{\rm F})/2}.$ (42)
This is an attractive interaction and its importance was well appreciated long ago by the systematic and unbiased survey of the electron-electron interaction in the problem of superconductivity in the electron gas [135]. Because the relevant interaction $V_{\rm ex}$ is attractive, we can understand why the excitron energy $\xi_{\bm{k}}$ is lower than ${\tilde{\varepsilon}}_{\bm{k}}$.
[…]
Figure 16: Overall structure of the renormalization function $Z(K)$ in (a) and that of the correlation part in the self-energy $\chi_{c}(K)$ in (b) for the electron gas characterized by the same parameters as those in Fig. 13.
After much trial and error, we come to realize that the mathematical feature is better seized in terms of $\Sigma({\bm{k}},i\omega_{n})$ rather than $\Sigma^{R}({\bm{k}},\omega\!+\!i\gamma)$, as long as only numerically obtained data are available to us at the present stage. Therefore, let us draw $Z(K)$ the renormalization function and $\chi_{c}(K)$ the correlation contribution to $\chi(K)$ in Fig. 16(a) and (b), respectively, as a function of $k$ in a wide range from 00 to $4k_{\rm F}$ [and even up to $10k_{\rm F}$ for $\chi_{c}(K)$] with changing Matsubara frequency $\omega_{n}$ also in a very wide range, i.e., for $n$ from $1$ ($\omega_{1}\approx 0.0006\varepsilon_{\rm F}$) to $7\times 10^{4}$ ($\omega_{70000}\approx 80\varepsilon_{\rm F}$). where $\chi_{c}(K)$ is defined by
$\chi_{c}(K)\equiv\chi(K)-\chi_{x}(k),$ (43)
with the exchange part of the self-energy $\chi_{x}(k)$ which is independent of $\omega_{n}$ and calculated as
$\frac{\chi_{x}(k)}{\varepsilon_{\rm F}}=-\frac{2\alpha r_{s}}{\pi}\left[1+\frac{k_{\rm F}^{2}-k^{2}}{2k_{\rm F}k}\ln\left|\frac{k_{\rm F}+k}{k_{\rm F}-k}\right|\,\right].$ (44)
As we see, in the scale of these figures, both $Z(K)$ and $\chi_{c}(K)$ seem to behave quite normally in the whole $\{k,\omega_{n}\}$ space, in accordance with FLT. In fact, even in a much-enlarged scale with $\omega_{n}$ in the limited range of $n$ from $1$ to $100$, we hardly see noticeable variations, much less anomaly, in $\chi_{c}(K)$ even for $k$ in the vicinity of $k_{\rm F}$.
Figure 17: Enlarged view of $Z({\bm{k}},i\omega_{n})$ given in Fig. 16(a) for $k/k_{\rm F}$ in the range of $0.94-1.06$ and $\omega_{n}$ with $n\lesssim 300$ (or $\omega_{n}\lesssim 0.38\varepsilon_{\rm F})$.
In $Z(K)$, however, we find a conspicuous spike structure for $k$ in the vicinity of $k_{\rm F}$ (or $0.99\lesssim k/k_{\rm F}\lesssim 1.01$) and small $\omega_{n}$ (or $\omega_{n}/\varepsilon_{\rm F}\lesssim 0.03$). Actually, this anomalous behavior can also be faintly seen in Fig. 16(a), but it is very clearly found through an enlarged view of $Z(K)$, as given in Fig. 17. This kind of anomaly in $Z(K)$ is never expected in FLT, but here it emerges as a strong symptom of the possible breakdown of FLT. At the same time, it is found to be very localized in $\{k,\omega_{n}\}$ space as a characteristic feature.
Figure 18: Division of $Z(K)$ given in Fig. 17 into (a) the smoothed part $Z^{\rm smooth}(K)$ and (b) the singular part $Z^{\rm diff}(K)$.
Because of this locality, we are tempted to extract this anomalous structure from $Z(K)$ by taking the difference between $Z(K)$ and $Z^{\rm smooth}(K)$, the latter of which is a smoothed part of $Z(K)$ obtained by the cubic-spline interpolation along $k$-axis with the exclusion of the mesh points in the interval $(0.97k_{\rm F},1.03k_{\rm F})$ at each $\omega_{n}$ with $n\leq 50$. For $n>50$, $Z^{\rm smooth}(K)$ is tentatively taken as $Z(K)$ itself. In Fig. 18(a) and (b), both $Z^{\rm smooth}(K)$ and $Z^{\rm diff}(K)\ [\equiv\!Z(K)-Z^{\rm smooth}(K)]$ are drawn, respectively. As we see, $Z^{\rm smooth}(K)$ is rather smooth and has no abnormal structure in the whole $\{k,\omega_{n}\}$ space. On the other hand, $Z^{\rm diff}(K)$ has a distinctive particle-hole symmetric structure and its magnitude is not small only in the very limited region in $\{k,\omega_{n}\}$ space, or more concretely, for $|k-k_{\rm F}|\!\lesssim\!0.025k_{\rm F}$ and $\omega_{n}\!\lesssim\!0.05\varepsilon_{\rm F}$ (or $n\lesssim 40$).
Incidentally, a similar extraction procedure cannot be adopted to produce $\chi_{c}^{\rm smooth}(K)$ and $\chi_{c}^{\rm diff}(K)$ at this stage, mainly because $\chi_{c}(K)$ does not change much with $\omega_{n}$ in the small-$\omega_{n}$ region, making it difficult to clearly identify the anomalous structure in $\chi_{c}(K)$. In Sec. IV.3, however, we shall explain an alternative procedure to unambiguously define both $\tilde{\chi}_{c}^{\rm smooth}(K)$ and $\tilde{\chi}_{c}^{\rm diff}(K)$ from the division of $\chi_{c}(K)$ in perfectly consistent with the redefined division of $Z(K)$ into $\widetilde{Z}^{\rm smooth}(K)$ plus $\widetilde{Z}^{\rm diff}(Z)$.
IV.3 Branch-cut singularity
Fascinated by the mathematically beautiful mirror-symmetry in $Z^{\rm diff}(K)$ with respect to $k$ in reference to $k_{\rm F}$ at each $\omega_{n}$, we proceed to express it in an analytically closed form. The discussion on $\Sigma_{\rm ex}$ and $V_{\rm ex}$ in Sec. IV.1 indicates that we might be able to treat our problem in reference to 1D physics, because the virtual scattering processes of an electron specified by ${\bm{k}}$ on the Fermi surface, along with the multiple electron-hole pair excitations to bring about the excitron, occur predominantly in either parallel or antiparallel to the ${\bm{k}}$-direction. Now, in some particularly simple models in the 1D Tomonaga-Luttinger liquids [136, 137], an analytic form for $G(K)\!\equiv\!G_{\rm TL}(K)$ is exactly known as [138]
$G_{\rm TL}(K)=\frac{1}{\sqrt{i\omega_{n}-\varepsilon_{h}(k)}\sqrt{i\omega_{n}-\varepsilon_{s}(k)}},$ (45)
where $\varepsilon_{h}(k)$ and $\varepsilon_{s}(k)$ are, respectively, “holon” and “spinon” dispersion relations in 1D physics.
Inspired by this simple expression for $G_{\rm TL}(K)$ with possessing branch-cut singularities, we shall take a heuristic approach to developing an analytic expression for $Z^{\rm diff}(K)$ by starting with the redefined division of $\Sigma(K)$ as
$\Sigma(K)=\widetilde{\Sigma}_{\rm smooth}(K)+\widetilde{\Sigma}_{\rm diff}(K),$ (46)
with
$\widetilde{\Sigma}_{\rm smooth}(K)=$ $\left[1-\widetilde{Z}^{\rm smooth}(K)\right]i\omega_{n}$
$+\chi_{x}(k)+\tilde{\chi}_{c}^{\rm smooth}(K),$ (47a)
$\widetilde{\Sigma}_{\rm diff}(K)=$ $-\widetilde{Z}^{\rm diff}(K)i\omega_{n}+\tilde{\chi}_{c}^{\rm diff}(K).$ (47b)
Here, $\widetilde{\Sigma}_{\rm smooth}(K)$ is regarded as such a self-energy in the electron gas as has been considered in FLT and $\widetilde{\Sigma}_{\rm diff}(K)$ is supposed to accurately take account of the anomalous structure in $\Sigma(K)$ by the assumption of the following analytic form:
$\widetilde{\Sigma}_{\rm diff}(K)=$ $-\sqrt{i\omega_{n}-\xi_{\bm{k}}}\sqrt{i\omega_{n}-\xi^{*}_{\bm{k}}(i\omega_{n})}$
$+\frac{\sqrt{\alpha_{\infty}(i\omega_{n})}}{2}(i\omega_{n}-\xi_{\bm{k}})$
$+\frac{1}{2\sqrt{\alpha_{\infty}(i\omega_{n})}}\left[i\omega_{n}-\xi^{*}_{\bm{k}}(i\omega_{n})\right],$ (48)
where $\xi_{\bm{k}}$ is chosen as the excitron dispersion relation given in Fig. 8 and $\xi^{*}_{\bm{k}}(i\omega_{n})$ is defined by $\xi^{*}_{\bm{k}}(i\omega_{n})\equiv\alpha_{\bm{k}}(i\omega_{n})\xi_{\bm{k}}$ with the introduction of a parameter $\alpha_{\bm{k}}(i\omega_{n})$. As we shall see, in the limit of $|\xi_{\bm{k}}/\omega_{n}|\gg 1$, $\alpha_{\bm{k}}(i\omega_{n})$ becomes independent of ${\bm{k}}$ and the value in this limit is written as $\alpha_{\infty}(i\omega_{n})$, a parameter involved in Eq. (48).
Physically, $\xi_{\bm{k}}$ is supposed to represent a “collective-charge” excitation (or a holon-like excitation in the terminology of 1D physics) and thus it is considered as the excitron dispersion. On the other hand, we presume that $\xi^{*}_{\bm{k}}(i\omega_{n})$ corresponds to a spinon-like excitation whose dispersion relation is not much different from the quasiparticle dispersion ${\tilde{\varepsilon}}_{\bm{k}}$ in the electron gas, leading us to the condition that $\alpha_{\bm{k}}(i\omega_{n})$ should not be less than unity due to the fact that $\xi_{\bm{k}}/\tilde{\varepsilon}_{\bm{k}}<1$ in the present case.
In this theoretical framework, $\alpha_{\bm{k}}(i\omega_{n})$ is a single free parameter and plays an important role in $\widetilde{\Sigma}_{\rm diff}(K)$; if $\alpha_{\bm{k}}(i\omega_{n})$ is taken to be unity, $\widetilde{\Sigma}_{\rm diff}(K)$ vanishes completely, implying that the strength of the anomaly is determined only by the degree of deviation of $\alpha_{\bm{k}}(i\omega_{n})$ from unity. We also note that the $\omega_{n}$-dependence in $\alpha_{\bm{k}}(i\omega_{n})$ seems to be indispensable in describing a branch-cut singularity in not exactly 1D systems, as opposed to purely 1D systems in which such dependence is absent as seen in Eq. (45). Effects of the motions tangential to the 1D axis on the branch-cut singularity are assumed, to a large extent, to be effectively included by this $\omega_{n}$-dependence.
In view of the importance of $\alpha_{\bm{k}}(i\omega_{n})$, we make a rather detailed explanation of the procedure to determine it by using the numerically-obtained data of $Z^{\rm diff}(K)$. From Eq. (48), we can write $\widetilde{Z}^{\rm diff}(K)$ and $\tilde{\chi}_{c}^{\rm diff}(K)$ as
$\widetilde{Z}^{\rm diff}(K)=$ $[S_{-}(K)S^{*}_{+}(K)+S_{+}(K)S^{*}_{-}(K)]/\omega_{n}$
$-\frac{\sqrt{\alpha_{\infty}(i\omega_{n})}}{2}\left[1+\frac{1}{\alpha_{\infty}(i\omega_{n})}\right],$ (49a)
$\tilde{\chi}_{c}^{\rm diff}(K)=$ $S_{+}(K)S^{*}_{+}(K)-S_{-}(K)S^{*}_{-}(K)$
$-\frac{\sqrt{\alpha_{\infty}(i\omega_{n})}}{2}\left[\xi_{\bm{k}}+\frac{\xi^{*}_{\bm{k}}(i\omega_{n})}{\alpha_{\infty}(i\omega_{n})}\right],$ (49b)
respectively, where $S_{\pm}(K)$ is defined as
$S_{\pm}(K)\equiv\sqrt{\frac{\sqrt{\omega_{n}^{2}+\xi_{\bm{k}}^{2}}\pm\xi_{\bm{k}}}{2}},$ (50)
and $S^{*}_{\pm}(K)$ is defined analogously with the replacement of $\xi_{\bm{k}}$ with $\xi^{*}_{\bm{k}}(i\omega_{n})$. From Eqs. (49a) and (49b), together with Eq. (50), it is easy to see that both $\widetilde{Z}^{\rm diff}(K)$ and $\tilde{\chi}_{c}^{\rm diff}(K)$ vanish in the limit of $|\xi_{\bm{k}}/\omega_{n}|\gg 1$.
Figure 19: Overall structure of the parameter $\alpha_{\bm{k}}(i\omega_{n})$ as a function of $k\!=\!|{\bm{k}}|$ for $\omega_{n}$ in the range of $n\!=\!1\!-\!50$ in (a), its asymptotic form in the limit $|\xi_{\bm{k}}|\!\gg\!|\omega_{n}|$, $\alpha_{\infty}(i\omega_{n})$, with the increase of $n$ in (b), and the parameter $\beta_{\bm{k}}$ related to the excitron dispersion relation as a function of $k\!=\!|{\bm{k}}|$ in (c).
Now, because of $\xi_{\bm{k}}\!=\!\xi^{*}_{\bm{k}}(i\omega_{n})\!=\!0$ at the fermi level, we obtain $\widetilde{Z}^{\rm diff}({\bm{k}}_{\rm F},i\omega_{n})\!=\!1\!-\![\sqrt{\alpha_{\infty}(i\omega_{n})}\!+\!1/\sqrt{\alpha_{\infty}(i\omega_{n})}]/2$. Then, By equating this value of $\widetilde{Z}^{\rm diff}({\bm{k}}_{\rm F},i\omega_{n})$ to the numerically obtained $Z^{\rm diff}({\bm{k}}_{\rm F},i\omega_{n})$, we can determine $\alpha_{\infty}(i\omega_{n})$ as
$\alpha_{\infty}(i\omega_{n})=\left(Z_{0}+\sqrt{Z_{0}^{2}-1}\right)^{2},$ (51)
with $Z_{0}\!\equiv\!1\!-\!Z^{\rm diff}({\bm{k}}_{\rm F},i\omega_{n})$. This value of $\alpha_{\infty}(i\omega_{n})$ is chosen under the condition of $\alpha_{\infty}(i\omega_{n})\geq 1$ and the obtained result is shown in Fig. 19(b) as a function of $n$. Actually, because $Z^{\rm diff}({\bm{k}}_{\rm F},i\omega_{n})$ is available only for $n\leq 50$, $\alpha_{\infty}(i\omega_{n})$ for $n>50$ is not determined by Eq. (51) but by $\alpha_{\infty}(i\omega_{n})\!=\!1\!+\!\delta\alpha_{\infty}(i\omega_{n})$ with the extrapolation of the data $\{\delta\alpha_{\infty}(i\omega_{n})\}_{n=1,\cdots,50}$ under the assumption of the power-law decay of $\delta\alpha_{\infty}(i\omega_{n})\ [\equiv\alpha_{\infty}(i\omega_{n})-1]$ with the increase of $n$ from $50$.
Once the data of $\{\alpha_{\infty}(i\omega_{n})\}_{n=1,2,3,\cdots}$ are known, we can determine $\alpha_{\bm{k}}(i\omega_{n})$ at each ${\bm{k}}$ by accurately solving the equation of $\widetilde{Z}^{\rm diff}(K)=Z^{\rm diff}(K)$ with the use of the Newton-Raphson method [139] in which we employ the partial derivative of $\widetilde{Z}^{\rm diff}(K)$ with respect to $\alpha_{\bm{k}}(i\omega_{n})$ as
$\frac{\partial\widetilde{Z}^{\rm diff}(K)}{\partial\alpha_{\bm{k}}(i\omega_{n})}=$ $\frac{\xi_{\bm{k}}}{2\omega_{n}\sqrt{\omega_{n}^{2}+{\xi^{*}_{\bm{k}}(i\omega_{n})}^{2}}}$
×$\times[S_{-}(K)S^{*}_{+}(K)-S_{+}(K)S^{*}_{-}(K)],$ (52)
and choose $\alpha_{\infty}(i\omega_{n})$ as an initial input in the iterative solution for $n\leq 50$. For $n>50$, $\alpha_{\bm{k}}(i\omega_{n})$ is determined by an extrapolation method similar to that for $\alpha_{\infty}(i\omega_{n})$. The obtained result of $\alpha_{\bm{k}}(i\omega_{n})$ for $n\leq 50$ is given in Fig. 19(a) from which we see that for $|k\!-\!k_{\rm F}|\gtrsim 0.02k_{\rm F}$ in the important range of $\omega_{n}$ with $n\lesssim 30$, $\alpha_{\bm{k}}(i\omega_{n})$ is independent of ${\bm{k}}$ and essentially the same as $\alpha_{\infty}(i\omega_{n})$. Even for $n\gtrsim 30$, we see that $\alpha_{\bm{k}}(i\omega_{n})\approx\alpha_{\infty}(i\omega_{n})$ for $|k\!-\!k_{\rm F}|\gtrsim 0.05k_{\rm F}$, revealing that $\alpha_{\infty}(i\omega_{n})$ is the most important parameter to describe $\widetilde{\Sigma}_{\rm diff}(K)$. In Fig. 19(c), the parameter $\beta_{\bm{k}}$ defined by $\beta_{\bm{k}}\equiv\xi_{\bm{k}}/\varepsilon_{\bm{k}}$ is also shown. Note that $\beta_{\bm{k}}$ at $|{\bm{k}}|\!=\!k_{\rm F}$ is nothing but $v_{\rm excitron}/v_{\rm F}$. The deviation of $\beta_{\bm{k}}$ from $\beta_{{\bm{k}}_{\rm F}}$ represents the degree of the departure from the linear dispersion in $\xi_{\bm{k}}$.
Figure 20: Singular part $\tilde{\chi}_{c}(K)$ in $\chi_{c}(K)$ as a function of $k=|{\bm{k}}|$ for small $\omega_{n}$ with $n\!=\!1\!-\!50$.
Having completely specified the parameter $\alpha_{\bm{k}}(i\omega_{n})$ in the whole $\{{\bm{k}},\omega_{n}\}$ space, we can employ Eqs. (49a) and (49b) to calculate both $\widetilde{Z}^{\rm diff}(K)$ and $\tilde{\chi}_{c}^{\rm diff}(K)$, with which we can also determine $\widetilde{Z}^{\rm smooth}(K)\ [\equiv\!Z(K)\!-\!\widetilde{Z}^{\rm diff}(K)]$ and $\tilde{\chi}_{c}^{\rm smooth}(K)\ [\equiv\!\chi_{c}(K)\!-\!\tilde{\chi}_{c}^{\rm diff}(K)]$ unambiguously. The obtained $\widetilde{Z}^{\rm diff}(K)$ and $\widetilde{Z}^{\rm smooth}(K)$ are, respectively, found to be virtually the same as $Z^{\rm diff}(K)$ and $Z^{\rm smooth}(K)$ given in Fig. 18 and thus we suppress to show those redefined functions here.
In Fig. 20, $\tilde{\chi}_{c}^{\rm diff}(K)$ is drawn to show that its behavior is qualitatively different from that of $\widetilde{Z}^{\rm diff}(K)$; as opposed to the locality of $Z^{\rm diff}(K)$ or $\widetilde{Z}^{\rm diff}(K)$, $\tilde{\chi}_{c}^{\rm diff}(K)$ is considerably extended in the $k$-axis. We also note that its magnitude is quite small, of the order of $0.001\varepsilon_{\rm F}$, compared with that of $\chi_{c}(K)$ of the order of $\varepsilon_{\rm F}$, making $\tilde{\chi}_{c}^{\rm smooth}(K)$ virtually the same as $\chi_{c}(K)$. Those features specific to $\tilde{\chi}_{c}^{\rm diff}(K)$ are the reasons why we could not find any anomalous behavior in $\chi_{c}(K)$ in the first place. Asymmetry with respect to $k$ in reference to $k_{\rm F}$ in $\tilde{\chi}_{c}^{\rm diff}(K)$ is another interesting point to note.
[…]
Figure 21: Comparison of $A({\bm{k}},\omega)$ with $A_{\rm smooth}({\bm{k}},\omega)$ the one-particle spectral function corresponding to the smoothed self-energy is made in (a) for the case of $k\!=\!|{\bm{k}}|\!=\!0.95k_{\rm F}$. The singular part of the retarded self-energy and the analytically continued ${\tilde{\alpha}}(\omega)$ from $\alpha_{\infty}(i\omega_{n})$ are, respectively, shown in (b) and (c).
At first glance, one might think that we can easily make an analytic continuation of $\widetilde{\Sigma}_{\rm diff}(K)$ to $\widetilde{\Sigma}_{\rm diff}^{R}({\bm{k}},\omega\!+\!i\gamma)$ by just changing $i\omega_{n}$ into $\omega+i\gamma$ in Eq. (48), but actually it is not so simple due to the presence of $\alpha_{\bm{k}}(i\omega_{n})$; the analyticity property of $\alpha_{\bm{k}}(i\omega_{n})$ is not precisely known. Thus, we employ the usual analytic continuation method (or Padé approximants) to obtain $\widetilde{\Sigma}_{\rm smooth}^{R}({\bm{k}},\omega\!+\!i\gamma)$ from $\widetilde{\Sigma}_{\rm smooth}(K)$ defined in Eq. (47a) and then we determine $\widetilde{\Sigma}_{\rm diff}^{R}({\bm{k}},\omega\!+\!i\gamma)$ as the difference between $\Sigma^{R}({\bm{k}},\omega\!+\!i\gamma)$ and $\widetilde{\Sigma}_{\rm smooth}^{R}({\bm{k}},\omega\!+\!i\gamma)$, the former of which is already obtained from $\Sigma(K)$. The result of $\widetilde{\Sigma}_{\rm diff}^{R}({\bm{k}},\omega\!+\!i\gamma)$ in the low-energy region is drawn in Fig. 21(b) which is perfectly consistent with the anomalous structure of $\Sigma^{R}({\bm{k}},\omega\!+\!i\gamma)$ around the excitron position in the inset in Fig. 13. As related to $\widetilde{\Sigma}_{\rm smooth}^{R}({\bm{k}},\omega\!+\!i\gamma)$, the one-particle spectral function $A_{\rm smooth}({\bm{k}},\omega)$ is given in Fig. 21(a) by the black dotted curve in comparison with $A({\bm{k}},\omega)$ indicated by the red solid curve, where $A_{\rm smooth}({\bm{k}},\omega)$ is defined by
$A_{\rm smooth}({\bm{k}},\omega)=-\frac{1}{\pi}{\rm Im}\,G_{\rm smooth}^{R}({\bm{k}},\omega\!+\!i\gamma)$ (53)
with
${G_{\rm smooth}^{R}({\bm{k}},\omega\!+\!i\gamma)}^{-1}=$ $\,\omega\!+\!i\gamma\!+\!\mu_{x}\!+\!\mu_{c}^{\rm smooth}\!-\!\varepsilon_{\bm{k}}$
$-\widetilde{\Sigma}_{\rm smooth}^{R}({\bm{k}},\omega\!+\!i\gamma),$ (54)
where $\mu_{c}^{\rm smooth}$ is the correlation contribution to the chemical potential as calculated through $\widetilde{\Sigma}_{\rm smooth}(K)$. Actually, its difference from $\mu_{c}$ is negligibly small. There exists no signiture of the excitron in $A_{\rm smooth}({\bm{k}},\omega)$ as it should be in the case of FLT.
In order to investigate the function analytically continued from $\alpha_{\bm{k}}(i\omega_{n})$, we write down $\Sigma^{R}({\bm{k}},\omega\!+\!i\gamma)$ in reference to Eq. (48) as
$\widetilde{\Sigma}_{\rm diff}^{R}({\bm{k}},\omega\!+\!i\gamma)=$ $-\sqrt{\omega\!+\!i\gamma-\xi_{\bm{k}}}\sqrt{\omega\!+\!i\gamma-{\tilde{\alpha}}(\omega)\xi_{\bm{k}}}$
$+\frac{\sqrt{{\tilde{\alpha}}(\omega)}}{2}(\omega\!+\!i\gamma-\xi_{\bm{k}})$
$+\frac{1}{2\sqrt{{\tilde{\alpha}}(\omega)}}\left[\omega\!+\!i\gamma-{\tilde{\alpha}}(\omega)\xi_{\bm{k}}\right],$ (55)
where ${\tilde{\alpha}}(\omega)$ is introduced as the analytically continued function from $\alpha_{\infty}(i\omega_{n})$, but because $\alpha_{\bm{k}}(i\omega_{n})$ is well approximated by $\alpha_{\infty}(i\omega_{n})$ at $|{\bm{k}}|\!=\!0.95k_{\rm F}$, it also represents the analytically continued one from $\alpha_{\bm{k}}(i\omega_{n})$. The branch cut in the square root $\sqrt{z}$ in Eq. (55) is taken along the negative real axis in complex-$z$ plane.
By comparing the result of $\widetilde{\Sigma}_{\rm diff}^{R}({\bm{k}},\omega\!+\!i\gamma)$ in Fig. 21(b) with that in Eq. (55), we can determine not rigorously correct but reasonably accurate values of ${\tilde{\alpha}}(\omega)$ for $\omega$ in the range from $-0.2\varepsilon_{\rm F}$ to 00 and the obtained results are given in Fig. 21(c), from which we can raise a couple of points to appreciate the importance of $\omega$-dependence in ${\tilde{\alpha}}(\omega)$: (i) Although ${\tilde{\alpha}}(\omega)$ is real at $\omega=0$ (more concretely, ${\tilde{\alpha}}(0)\!=\!2.247$ as given by an extrapolation of the data $\{\alpha_{\infty}(i\omega_{n})\}_{n=1,2,3,\cdots}$), it is generally complex with a negative imaginary part. Due to the existence of this imaginary part in ${\tilde{\alpha}}(\omega)$, $A({\bm{k}},\omega)$ is not characterized by such a square-root singularity, $1/\sqrt{\xi_{\rm k}-\omega}$, as typically seen in purely 1D Luttinger liquids [140, 141, 82, 142], but is well approximated by a Lorentzian-type function [143]. (ii) One might expect to see another anomaly at $\omega={\tilde{\alpha}}(\omega)\xi_{\bm{k}}$ originating from the second square root $\sqrt{\omega\!+\!i\gamma-{\tilde{\alpha}}(\omega)\xi_{\bm{k}}}$ in Eq. (55), but it does not seem to be the case, primarily because the deviation of ${\tilde{\alpha}}(\omega)$ from unity at the relevant $\omega$ is not large enough to provide a noticeable structure in $A({\bm{k}},\omega)$. The only effect from this contribution is found to make the quasiparticle peak position in $A({\bm{k}},\omega)$ shift slightly from that in $A_{\rm smooth}({\bm{k}},\omega)$ as seen in Fig. 21(a).
IV.5 $T$-dependence of the excitron liftime
In order to better understand the excitron, it is useful to determine its lifetime $\tau_{\rm excitron}$, especially, its dependence on $T$ in connection with the coherence nature of the excitron. In the conventional Fermi liquids, the quasiparticle lifetime $\tau_{*}$ is obtained by
$\tau_{*}^{-1}=2z^{*}|{\rm Im}\Sigma^{R}({\bm{k}}_{\rm F},0)|,$ (56)
and scales as $T^{2}$. Thus, on general grounds, one expects that ${\rm Im}\Sigma^{R}({\bm{k}_{\rm F}},\omega)$ also contains some relevant information on $\tau_{\rm excitron}^{-1}$, but because the excitron peak is hidden behind the quasiparticle peak at $k=k_{\rm F}$, it is not easy to extract information on the excitron from ${\rm Im}\Sigma^{R}({\bm{k}}_{\rm F},\omega)$ without removing the dominant quasiparticle contribution in the first place. According to the first-Matsubara-frequency rule [53], the quasiparticle part vanishes in the $T$-dependent ${\rm Im}\Sigma({\bm{k}}_{\rm F},i\pi T)$, implying the possibility that we may directly connect ${\rm Im}\Sigma({\bm{k}}_{\rm F},i\pi T)$ with $\tau_{\rm excitron}^{-1}$.
Figure 22: (a) Change of $\Sigma^{R}({\bm{k}},\omega\!+\!i\pi T)$ with $T$ in the range from $10^{-4}\varepsilon_{\rm F}$ to $8\!\times\!10^{-4}\varepsilon_{\rm F}$ at $k=0.99k_{\rm F}$, $k=k_{\rm F}$, and $1.01k_{\rm F}$ for $r_{s}=3.93$. (b) Temperature dependence of $A({\bm{k}},i\pi T)$ at $k=0.99k_{\rm F}$ and $1.01k_{\rm F}$ for $T$ in the same range as that in (a).
In these circumstances, we have examined the $T$-dependence of $\Sigma^{R}({\bm{k}},\omega\!+\!i\pi T)$ not only at $k\!=\!k_{\rm F}$ but also in the very vicinity of $k_{\rm F}$, namely, at $k\!=\!0.99k_{\rm F}$ and $k\!=\!1.01k_{\rm F}$ and the obtained results are shown in Fig. 22(a). The corresponding ones for $A({\rm k},\omega)$ are also given in Fig. 22(b) from which we estimate $\tau_{\rm excitron}^{-1}$ by measuring a full width at a half maximum (FWHM). Note that we cannot obtain this FWHM unambiguously, mainly because the excitron peak is located at the foot of the dominant quasiparticle peak. Thus, a rather large error bar is associated with this estimation, particularly for $T\!\gtrsim\!4\!\times\!10^{-4}\!\varepsilon_{\rm F}$. Those data for $\tau_{\rm excitron}^{-1}$ can be summarized in the form of $\tau_{\rm excitron}^{-1}\!\propto\!T^{\alpha}$ with $\alpha\!\approx\!1.0\!\pm\!0.2$. Similar analysis is also made in reference to $-{\rm Im}\Sigma^{R}({\bm{k}},i\pi T)$ to find $\alpha\!\approx\!0.6\!-\!0.9$. Incidentally, we find that the numerical data for ${\rm Im}\Sigma^{R}({\bm{k}},i\pi T)$ are very accurately expressed by
${\rm Im}\Sigma^{R}((1\pm 0.01){\bm{k}}_{\rm F},i\pi T)=-(2.22\mp 0.02)T^{1.00},$ (57)
at $k\!=\!(1\!\pm\!0.01)k_{\rm F}$, while at $k\!=\!k_{\rm F}$,
${\rm Im}\Sigma^{R}({\bm{k}}_{\rm F},i\pi T)=-0.924\varepsilon_{\rm F}(T/\varepsilon_{\rm F})^{0.91},$ (58)
for $T$ less than $8\!\times\!10^{-4}\!\varepsilon_{\rm F}$. If we assume that $\tau_{\rm excitron}^{-1}$ is directly proportional to ${\rm Im}\Sigma^{R}({\bm{k}}_{\rm F},i\pi T)$, we obtain another estimate of $\alpha$ as $\alpha\!=\!0.91$. In this way, $\alpha$ seems to be close to unity which is exactly the value evaluated in 1D Luttinger liquids [144]. The result of $\alpha=1$ is also expected in the marginal Fermi liquids. In the present case, however, it is likely that $\alpha$ is less than unity, though a definite value of $\alpha$ should be determined by some other analytic method in the future.
Because of $\alpha<1$, the excitron is not a coherent but an incoherent excitation. Thus, it should formally be included in the incoherent background in Eq. (1), but it still provides a clear peak structure in $A({\rm k},\omega)$ contrary to the general belief that the incoherent background will be represented by a smooth function in the Fermi liquids.
Intuitively, we can think of the following: The diagram in Fig. 15(c) indicates that the excitron is an electron surrounded by multiple electron-hole pairs excited mostly in the longitudinal direction. In this sense, the excitron (or an electron-multiexciton complex in a shape elongated along the electron motion) may be regarded as a kind of polarons. In the case of polarons, however, the associated mediating modes are phonons which are coherent bosons in the whole crystal. On the other hand, the excitron is associated with the incipient excitonic mode, an incoherent boson mode which is damped in a short distance. Thus, it is very reasonable to reach the conclusion that the excitron is an incoherent excitation.
V Conclusion and Discussion
In this paper, we have developed a feasible nonperturbative scheme to accurately determine $\Sigma(K)$ through a fully self-consistent iterative calculation with rigorously satisfying the Ward identity and the total momentum conservation law, while fulfilling all other known conservation laws, sum rules, and correct asymptotic behaviors in $G(K)$, $\Gamma(K,K\!+\!Q)$, $\Pi(Q)$, and $n({\bm{k}})$. The scheme has been successfully implemented in the 3D homogeneous electron gas for the range of $r_{s}$ corresponding to all simple metals at $T$ down to $10^{-4}\varepsilon_{\rm F}$ with tiny mesh as small as $10^{-4}k_{\rm F}$ near the Fermi surface in ${\bm{k}}$ space. Our results on $n({\bm{k}})$, the quasiparticle renormalization factor $z^{*}$, and the quasiparticle effective mass $m^{*}$, all of which are the long-standing challenges in the electron gas, are in very good agreement with the recent data given by quantum Monte Carlo simulations and available experiments, confirming that our present scheme actually provides sufficiently accurate results of $\Sigma(K)$.
By analytic continuation onto the real $\omega$ axis through Padé approximants, $G(K)$ is transformed into $G^{R}({\bm{k}},\omega)$, from which we obtain $A({\bm{k}},\omega)$ exhibiting a new sharp low-energy peak for $|{\bm{k}}|$ not just at $k_{\rm F}$ but in its vicinity, in addition to the dominant quasiparticle peak as well as high-energy one- and two-plasmon satellites. The appearance of two-plasmon satellites without resort to the ad hoc combination of the $GW$ approximation with a cumulant expansion is a notable theoretical achievement, but the most important issue is the discovery of the new low-energy peak that emerges for all simple-metal densities at $T\lesssim 10^{-3}\varepsilon_{\rm F}$. Its origin is attributed to the excitonic attraction arising from the multiple excitations of tightly-bound electron-hole pairs in $\Pi({\bm{q}},\Omega)$ for $|{\bm{q}}|\approx 2k_{\rm F}$ and $|\Omega|\ll\varepsilon_{\rm F}$, suggesting that it should be dubbed “excitron”. This excitonic scattering process occurs only in very restricted angles along the longitudinal direction, which motivates us to characterize the excitron as a branch-cut singularity in analogy with 1D physics. From a viewpoint of QCP physics, the excitron is also regarded as an anomaly induced by quantum fluctuations of the incipient excitonic mode around the quantum-critical CDW transition. In either way, this anomalous low-energy phenomenon poses an interesting question as to the validity of the Landau’s hypothesis on the one-to-one correspondence of low-energy excitations between a free Fermi gas and an interacting normal Fermi liquid. Taken together, our results indicate that non-Fermi liquid physics may already play a role in the description of simple metals at sufficiently low temperatures.
Four comments are in order:
(i) Since we start with the rigorous equation to determine $\Sigma(K)$ in Eq. (23) in which $W(Q)$ is taken as an accurately known quantity, the vertex function $\Gamma(K,K\!+\!Q)$ is the only unknown quantity. Thus, we have examined various forms for $\Gamma(K,K^{\prime})$ in our theoretical framework, looking for a necessary and sufficient condition for the appearance of excitron. As a result, we come to know that the excitron appears, if and only if we include either $\overline{\Gamma}_{\rm WI}(K,K^{\prime})$ in Eq. (37d) or $\Gamma_{\rm WI}(K,K^{\prime})$ without $\eta_{1}(Q)$ in Eq. (29) in the definition of $\Gamma(K,K^{\prime})$. We can easily understand the necessity of this kind of the Ward-identity-related vertex part, because this is the crux to make our scheme nonperturbative; remember that we need to go beyond simple perturbation expansion from $G_{0}(K)$ for describing a situation intimately connected with the breakdown of the Fermi-liquid theory. Incidentally, the quantitative details of excitron, such as the peak position $\xi_{\bm{k}}$, the peak height, and the peak width, depend on the choice of $\Gamma(K,K^{\prime})$ by not negligible amounts. Therefore, more useful information on excitron is needed in the future to further improve on $\Gamma(K,K^{\prime})$.
(ii) In the simple-metal density region, the strength of the excitron peak is found to be so weak that its existence will not be detected by bulk measurements such as electric conductivity and specific heat. In ARPES experiments, on the other hand, it will be detected, if the energy resolution is much smaller, of the order of 1 meV or less, than those in preceding experiments. So far, experimental equipment with the energy resolution of the order of 1 meV or less [145] has not been applied to simple metals, but if they are actually applied with due attention to the possible appearance of excitron, it will be very exciting to see the experimental results from a perspective of fundamental physics. Detection of excitron by ARPES is also very important from a viewpoint of further developments of our theoretical framework, as mentioned in the previous paragraph.
(iii) By making the electron density lower than those of simple metals to approach the CDW transition, we can expect a more interesting situation in which the effects of excitron become so strong that FLT apparently breaks down, leading to the emergence of NFL. With this expectation in mind, we are now trying to obtain a fully self-consistent solution for such low densities (or $r_{s}>6$), but at present it takes too many iterative steps to obtain a completely convergent result of $\Sigma(K)$ for $r_{s}>8$. We shall report our efforts in this direction in the near future.
(iv) In Sec. IV, we have not taken a mathematically rigorous but a heuristic approach to the analysis of the excitron. Admittedly, it would be better to derive the branch-cut singularity in $\Sigma(K)$ analytically by explicitly including the effect of $V_{\rm ex}(K,K^{\prime};Q)$ defined in Fig. 15(a), but we have to understand that this is a very difficult task. In fact, this problem of accurately treating the local charge fluctuations induced by correlated multiple electron-hole pair excitations is as difficult as that of the local spin fluctuations in the heavy-Fermion superconductors [146, 147, 25, 148, 149, 150, 151] and high-$T_{c}$ cuprate superconductors [152, 153, 154, 155, 16, 17, 18, 21], suggesting that we should leave this problem for future analysis. To put it the other way around, our present approach to treating $\Sigma(K)$ as a whole by imposing various conservation laws and sum rules may provide a new route to the solution of local spin fluctuation problems in those strongly-correlated materials. We would expect a new development from this perspective in those hot fields.
Acknowledgements.
The author would like to thank Kazuhiro Matsuda for useful discussions at the very early stage of this work. He is also grateful to Hiroyuki Yata and Naoki Kawashima at ISSP, the University of Tokyo for the supports for long years to maintain the cluster machines to efficiently perform the computations reported in this paper.
Appendix A Matsubara sum
The Matsubara sum of a given function $f(i\omega_{n})$ with $\omega_{n}=\pi T(2n-1)$ for an integer $n$ is calculated numerically in the following way:
$T\sum_{\omega_{n}}f(i\omega_{n})$ $=T\sum_{n=1}^{\infty}F(\omega_{n})$
$=T\sum_{n=1}^{N}F(\omega_{n})+\frac{T}{12}[F(\omega_{N})+5F(\omega_{N+1})]$
$\hskip 68.28644pt+\int_{\omega_{N+1}}^{\infty}\frac{dx}{2\pi}\,F(x),$ (59)
where $F(\omega_{n})\equiv f(i\omega_{n})+f(-i\omega_{n})$ and we increase $N$ from $N\approx 10$ until a convergent result is obtained; in most cases, $N$ as small as $10$ is already good enough, but it is safe to choose $N=100$. We can derive Eq. (59) from the Euler-Maclaurin formula [156]:
[…]
$\ +\frac{1}{2}F(b)\Bigr{]}-\frac{B_{2}}{2!}h^{2}F^{\prime}(x)\Bigr{|}_{a}^{b}-\cdots,$ (60)
with the Bernoulli number $B_{2}=1/6$. By taking $h\!=\!2\pi T$, $a\!=\!\omega_{N+1}$, $b\to\infty$, and $F^{\prime}(a)\!=\![F(a)\!-\!F(a\!-\!h)]/h$, we can easily arrive at Eq. (59). The relative error incurred in cutting off the series in Eq. (60) at the level of $F^{\prime}(x)$ is of the order of $T^{4}$, negligibly small for sufficiently low $T$.
Appendix B Modification of the momentum distribution function
We take the following strategy to improve on $n({\bm{k}})$:
(i) Obtain $n({\bm{k}})\ [=\!n(x)$ with $x\!=\!|{\bm{k}}|/k_{\rm F}]$ through Eq. (10). (ii) Determine $n_{0}\!\equiv\!n(0)$ and $n_{\pm}\!\equiv\!n(1\!\pm\!0^{+})$. (iii) With the use of these $n_{0}$ and $n_{\pm}$, obtain $n_{\rm IGZ}(x)$ in the parametrization scheme described in Ref. [54]. Note that “IGZ” stands for “improved Gori-Giorgi and Ziesche” [157]. (iv) Because $n_{\rm IGZ}(x)$ is almost the same as $n(x)$, construct a corrected function $n_{c}(x)$ which changes smoothly from $n(x)$ for $x\!\lesssim\!1.1$ to $n_{\rm IGZ}(x)$ for $x\!\gtrsim\!2.0$. Concretely, we define $n_{c}(x)$ as $n_{c}(x)\!\equiv\!n(x)$ for $x\leq x_{c}$ and $n_{c}(x)\!\equiv\!n_{\rm IGZ}(x)\!+\!\Delta n(x)$ for $x\!>\!x_{c}$ with choosing an appropriate $x_{c}$ in the region of $2.01.1\!<\!x_{c}\!<\!$ 2.0. The small additional term $\Delta n(x)$ decreases exponentially as $x$ increases. At $x\!=\!x_{c}$, $\Delta n(x)$ is so determined as to make $n_{c}(x)$ smoothly connected to $n(x)$ up to second derivative. It is also tuned to satisfy the three sum rules for the momentum distribution function as accurately as possible.
[…]
The behavior of $\Gamma(K,K\!+\!Q)$ at $K\!=\!K_{\rm F}$ in the limit of $Q\to Q_{0}$ is exactly known; in the $\omega$-limit, it approaches $\Gamma^{\omega}(K_{\rm F},K_{\rm F})$, given by
$\Gamma^{\omega}(K_{\rm F},K_{\rm F})=\left.\frac{\partial G^{-1}(K)}{\partial(i\omega_{n})}\right|_{K=K_{\rm F}}=Z(K_{\rm F}),$ (61)
as a direct consequence of WI, while in the $q$-limit, it approaches $\Gamma^{q}(K_{\rm F},K_{\rm F})$, given by
$\Gamma^{q}(K_{\rm F},K_{\rm F})=\left.\frac{\partial G^{-1}(K)}{\partial\mu}\right|_{K=K_{\rm F}}=\frac{\kappa}{\kappa_{\rm F}}\frac{\partial E(K_{\rm F})}{\partial\varepsilon_{{\bm{k}}_{\rm F}}},$ (62)
as one can convince oneself by considering the one-to-one correspondence of each Feynman diagram representing $\Gamma(K_{\rm F},K_{\rm F})$ with the one obtained by the differentiation of an arbitrary $G$ line in each Feynman diagram for $\Sigma$ with respect to $\mu$ [158, 159, 79]. If we use the expression in Eq. (25) as it is, the above limiting behavior is automatically satisfied, but in arriving at Eq. (33), we have introduced a few approximations and simplifications which may deteriorate this favorable feature. In fact, $\Gamma(K,K\!+\!Q)$ in Eq. (33) reduces to $\Gamma^{\omega}(K_{\rm F},K_{\rm F})/\eta_{1}(Q_{0})$ in the $\omega$-limit and to $\Gamma^{q}(K_{\rm F},K_{\rm F})[1/I_{0}+3\zeta_{3}(Q_{0})]$ in the $q$-limit, compelling us to impose the following constraints; $\eta_{1}(Q_{0})$ in the $\omega$-limit, $\eta_{1}^{\omega}\!=\!1$ and $\zeta_{3}(Q_{0})$ in the $q$-limit, $\zeta_{3}^{q}\!=\!(I_{0}\!-\!1)/(3I_{0})$. Note that $\eta_{1}(Q_{0})$ in the $q$-limit should be equal to $\eta_{1}^{q}\!\equiv\!{\tilde{\eta}}_{1}(K_{\rm F})\!=\!(\partial E(K_{\rm F})/\partial\varepsilon_{{\bm{k}}_{\rm F}})/Z(K_{\rm F})$ from the very definition of $\eta_{1}(Q)$.
Taking account of those constraints as well as the basic feature that $\Gamma(K,K\!+\!Q)$ should rapidly approach unity for either $K$ or $K\!+\!Q$ far away from $K_{\rm F}$, we have chosen $\zeta_{3}(Q)$ in Eq. (34) and $\eta_{1}(Q)$ in the following form:
$\eta_{1}(Q)=\frac{1}{[\eta_{a}^{-1}(q)-1]\eta_{b}(i\omega_{q}/2)\eta_{c}(Q)+1},$ (63)
where, $\eta_{a}(q)$ is the angular average of ${\tilde{\eta}}_{1}({\bm{k}},0)$ with respect to the angle $\theta$ between ${\bm{k}}_{\rm F}$ and ${\bm{q}}$ in the definition of ${\bm{k}}\!\equiv\!{\bm{k}}_{\rm F}\!+\!{\bm{q}}/2$ (and thus $k^{2}\!=\!k_{\rm F}^{2}\!+\!k_{\rm F}q\cos\theta\!+\!q^{2}/4)$, given by
$\eta_{a}(q)$ $=\langle{\tilde{\eta}}_{1}({\bm{k}},0)\rangle\equiv\frac{\int_{0}^{\pi}\sin\theta d\theta\,{\tilde{\eta}}_{1}({\bm{k}},0)}{\int_{0}^{\pi}\sin\theta d\theta}$
$=\int_{|k_{\rm F}\!-\!q/2|}^{k_{\rm F}\!+\!q/2}\frac{kdk}{k_{\rm F}q}\,\frac{\partial E({\bm{k}},0)/\partial\varepsilon_{\bm{k}}}{Z({\bm{k}},0)}.$ (64)
The overall $\omega_{q}$-dependence is described by $\eta_{b}(i\omega_{n})$ with $i\omega_{n}\to i\omega_{q}/2$ and the functional $\eta_{b}(i\omega_{n})$ is defined as
$\eta_{b}(i\omega_{n})=\frac{Z({\bm{k}}_{\rm F},i\omega_{n})-1}{Z({\bm{k}}_{\rm F},0)-1},$ (65)
by considering the fact that the dominant $\omega_{n}$ dependence comes from $Z(K)$ in $\Sigma(K)$ for $K$ near $K_{\rm F}$. The conversion function from $\omega$- to $q$-limits, $\eta_{c}(Q)$, is taken as
$\eta_{c}(Q)=\frac{v_{\rm F}^{2}q^{2}}{v_{\rm F}^{2}q^{2}+3\omega_{q}^{2}},$ (66)
in reference to the conversion in $\Pi_{0}(Q)$ at $Q\to Q_{0}$, as shown in Eqs. (19) and (20).
Appendix D Extrapolation to static quantity
From a set of data $\{f_{1},f_{2},\cdots,f_{N}\}$ for an even function $f(i\omega_{n})$ at $n=1,\cdots,N$ with $\omega_{n}\!=\!\pi T(2n\!-\!1)$, we can estimate the static value $f(0)$ by the following extrapolation: First, we regard the data set as that of the size twice as large by considering $\{f_{N},f_{N-1},\cdots,f_{2},f_{1},f_{1},f_{2},\cdots,f_{n}\}$ given at $\{-\omega_{N},-\omega_{N-1},\cdots,-\omega_{2},-\omega_{1},\omega_{1},\omega_{2},\cdots,\omega_{N}\}$. Then, we apply a Lagrange’s polynomial interpolation formula to this enlarged data set to obtain the interpolation function $\tilde{f}(\omega)$ as
$\tilde{f}(\omega)=\sum_{n=1}^{N}f_{n}\prod_{j\neq n}^{N}\frac{\omega^{2}-\omega_{j}^{2}}{\omega_{n}^{2}-\omega_{j}^{2}}.$ (67)
By substituting $x=0$ in Eq. (67), we obtain $f(0)$ as
$f(0)\approx\tilde{f}(0)=\sum_{n=1}^{N}f_{n}\prod_{j\neq n}^{N}\frac{\omega_{j}^{2}}{\omega_{j}^{2}-\omega_{n}^{2}}.$ (68)
We can check the convergence of the result $f(0)$ by increasing $N$ from $N\approx 10$ to find that in all cases $N=30$ is large enough.
Appendix E Mesh points
The selected mesh points in $(k,\omega_{n})$ space, $\{k_{i},\omega_{j}\}$, are chosen in the following way: On $k$-axis, the first point $k_{1}$ is taken as $k_{1}\!=\!0.007k_{\rm F}$; for $0\!<\!k_{i}\!\leq\!2k_{\rm F}$ and $|k_{i}\!-\!k_{\rm F}|\!\gtrsim\!0.01k_{\rm F}$, $\Delta_{i}\ (\equiv\!k_{i+1}\!-\!k_{i})\sim|k_{i}\!-\!k_{\rm F}|/10$; for $k_{i}\sim k_{\rm F}$, $\Delta_{i}\sim 10^{-4}k_{\rm F}$; for $k_{i}\!>\!2k_{\rm F}$, $\Delta_{i}\sim k_{i}/10$; and the last point $k_{M}$ is taken as $120k_{\rm F}$ with $M=240$. Since the minimum value of $\{\Delta_{i}\}$ is as small as $10^{-4}k_{\rm F}$, the integrals in Eqs. (36a)-(36c) should be very accurately performed, i.e., up to at least 6 digits, in order to obtain significant difference between the results at adjacent points. To achieve this accuracy, we employ the double-exponential formula for numerical integration [160].
As for $\omega_{n}$-axis, we take account of only the positive side of the axis, because we consider only even functions; for $j=1-24$, $\omega_{j}=\pi T(2j-1)$; for $j=25-35$, $\omega_{j}=\pi T(4j-49)$; for $j=36-50$, $\omega_{j}=\pi T(8j-189)$; $\cdots$ ; and the last point $\omega_{N}$ is about $10^{7}\times\pi T$ with $N=310$. This set $\{\omega_{j}\}$ is used in the Padé approximants for the numerical analytic continuation [126] in which it is quite useful to calculate in quadruple precision.